### v3

# The Tantalizing Case of the Quasar J09505128 – I. Presentation of the Data and Detailed Exploration of the Binary Supermassive Black Hole Scenario

Abstract
Spectroscopic observations of the quasar J09505128 spanning 22 years reveal monotonic radial velocity variations in its broad H emission line. Moreover, the line profile becomes broader over time, necessitating careful measurements. We present robust H velocity shift measurements obtained via cross correlation, applied to both the full spectra and to isolated broad H components derived from spectral decomposition. We also examine the light curves for variability consistent with the spectroscopic trends. Using Lomb-Scargle periodogram analysis we find no significant periodic signal. We consider several interpretations for the observed changes, including a binary supermassive black hole, dust-cloud obscuration, outflows, a recoiling black hole, and a single perturbed, disk-like broad-line region. We disfavor all but the binary and perturbed broad-line region scenarios. The binary interpretation is the only one for which we can immediately compare a physical model to the available data. Thus, we incorporate radial velocity “jitter” to emulate typical quasar variability and fit the radial velocity curve with a Keplerian model to examine whether it can reproduce the observations. In this context, the available observations trace only a segment of the putative orbit. The fit yields a period of 33 years (observed frame) and an eccentricity of 0.65, with lower limits on the semi-major axis and black hole mass of pc and , respectively. Thus, J09505128 is a compelling supermassive binary candidate. The single, perturbed broad-line region interpretation remains viable but requires additional observations and modeling for further evaluation. Continued monitoring is, therefore, essential.
1 Introduction
1.1 Background and Motivation
Galaxy mergers are a frequently observed phenomenon, and given that most massive galaxies have central supermassive black holes (BHs; kormendy1995inward), supermassive black hole binaries (SBHBs) are thought to be an inevitable consequence of the merger process. begelman1980massive described how a SBHB evolves in a post-galaxy merger. The evolution begins with the stellar bulges and respective BHs of the two galaxies sinking toward the center of the merger remnant. This process is initially driven by dynamical friction (chandrasekhar1943dynamical), which acts as the dominant mechanism for angular momentum loss for 109 yr for typical masses and can bring the BH pair closer. At a separation of pc, the BHs form a gravitationally bound binary (e.g., bogdanovic2022electromagnetic). Eventually, the binary “hardens” as its orbital velocity becomes comparable to the velocity dispersion of the stars. At this point, stellar scattering takes over as the dominant mechanism that further reduces the binary separation. Stellar scattering is effective so long as the number of stars with the right orbits to interact with the binary can be replenished (milosavljevic2003long; yu2005evolution). If this reservoir of stars is depleted, the binary separation may slow its evolution or stall at a separation of pc (known as the “last parsec problem”). However, several models––ranging from angular momentum loss through interactions with a circumbinary accretion disk or gaseous cloud (e.g., armitage2002accretion; escala2004role; dotti_lisa2006; cuadra2009massive) to dynamical models that consider non-spherical and anisotropic stellar distributions (e.g., merritt2004chaotic; berczik2006efficient; khan2012mergers; khan2013supermassive; holley2015galaxy)––now suggest that there are a number of sufficiently efficient mechanisms for continued orbital decay. Once the binary separation reaches 10-2 pc, angular momentum loss via gravitational waves (GWs) becomes an efficient mechanism, leading the two BHs to coalesce within a Hubble time.
Unambiguous observational evidence for SBHBs has been confined to the early stages of their evolution. Galaxy mergers and widely separated (kilo-parsec and sub-kilo-parsec scale) BH pairs are observed (e.g., comerford2009inspiralling; comerford20091; liu2011dual; ellison2017discovery; hou2019active; foord2020second; goulding2019dual; shen2021dual; koss2023dual), but bound SBHBs have yet to be imaged and resolved directly. An exception is the radio-loud candidate CSO 0402+379, directly imaged through radio interferometry with a projected separation of 7 pc (rodriguez2006compact; bansal2017constraining). The recent evidence of a stochastic GW background by pulsar timing array (PTA) experiments is compatible with the low-frequency signal expected from a population of SBHBs at the last stages of their evolution (agazie2023nanograv; agazie2023binary; antoniadis2023second; reardon2023search; lee2023searching; miles2025meerkat), strongly suggesting that SBHBs must evolve to merger, emitting GWs in the process. This is promising for directly resolving nanohertz GWs (corresponding to orbital periods of tens of years) from individual SBHBs at sub-parsec separations. The upcoming Laser Interferometer Space Antenna (LISA; amaro2023astrophysics; colpi2024) is also expected to detect SBHB mergers, at higher frequencies corresponding to somewhat lower BH masses and later stages of SBHB evolution than those detected by the PTAs.
SBHBs have also been invoked to explain a number of galaxy properties, including mass deficits in the cores of elliptical galaxies (e.g., milosavljevic2002cores; merritt2006cores), helical or precessing jets (e.g., roos1993precessing; romero2000preccessing), and X-shaped radio sources (e.g., merritt2002xshaped; gopalkrishna2003xshaped). Finding unambiguous SBHBs at small separations ( 0.11 pc) would greatly enhance our understanding of their evolution and their effects on their host galaxies. Thus, indirect methods have been developed to search for their observational signatures. The two most commonly used methods, which we describe further below, rely on periodic photometric variability or regular spectroscopic (radial velocity) variability of the broad emission lines. In addition to these, a few other methods have been proposed that are appropriate for SBHBs at separations of – 1 pc: peculiar profiles and relative intensities of broad emission lines (e.g., montuori11; montuori12), proper motion measurements via infrared interferometry (e.g. dexter20), periodic modulation of the polarization of scattered light (e.g. dotti22), kinematic signatures in reverberation mapping signals (e.g., wang18), or blue-red asymmetric reverberation of the broad emission lines (e.g. dotti23).
The photometric method involves looking for periodic optical flux variability in the light curves of quasars (e.g., graham2015systematic; charisi2016population; liu2016systematic) under the assumption that either one or both BHs are accreting, and that the accretion rate is regulated on the binary orbital period (d2013accretion). Periodic oscillations can also occur due to relativistic Doppler boosting from an orbiting active BH (dorazio2023review). A prominent photometric SBHB candidate is OJ 287, whose long-term light curve displayed outbursts separated on a 12-year period (e.g., valtonen2008Nature; valtonen2012; komossa2023oj27_1; komossa2023oj287_2). The use of periodic photometric variations to identify binaries, especially when only a small number of apparent orbital cycles is observed, however, is subject to ambiguities (liu2018did; witt2022period; robnik2024period), since quasars exhibit stochastic variability that can appear periodic (vaughan2016false; liu2016systematic; barth_stern2018; elbadry2025arXiv).
The spectroscopic method looks for the displaced peaks of broad emission lines relative to the host galaxy rest frame in quasar spectra (e.g., gaskell1983quasars; gaskell1996evidence; gaskell1996supermassive)––this is the approach used in this paper. A single-peaked broad emission line that is significantly offset is a possible signature of bulk motion of a single active BH and its surrounding gas in a binary. The possibility that the double-peaked emitters of some quasars can arise from systems with two active BHs has been tested and rejected based on the lack of radial velocity variations that are indicative of orbital motion (eracleous1997rejection; liu2016doublepeak; doan2020improved). Rather, double-peaked emission lines are best explained as originating in a disk-like broad-line region around a single supermassive black hole (see, for example, eracleous2003radioloud; eracleous2009review, for summary of arguments). While the single-peaked offset line method is viable for SBHB searches, it suffers from the ambiguity that offsets of broad emission lines can also be caused by other effects, such as outflows from the accretion disk (e.g., popovic2012super) or large, non-axisymmetric perturbations of the broad-line region (BLR; e.g., gezari2007long; lewis2010long; storchi2017double; schimoia2017evolution). It is therefore necessary to have multiple observations over extended periods of time to evaluate SBHB candidates identified by this method.
In order to select SBHB candidates based on their displaced broad emission lines, one approach is to search for 1000 km s-1 velocity offsets of the broad H emission lines relative to the narrow emission lines in the spectra of quasars (gaskell1983quasars; tsalmantza2011systematic; eracleous2012large; decarli2013nature; liu2014constraining). Using this method, eracleous2012large selected 88 SBHB candidates which have since undergone extensive studies to characterize their nature such as flux variability, radial velocity changes, and implied population properties (runnoe2015large; runnoe2017large; pflueger2018likelihood; nguyen2020pulsar). runnoe2017large used observations through 2011 and found 29 candidates that displayed steady, and in some cases, statistically significant radial velocity variations. Among these, three objects displayed monotonic changes consistent with the expected behavior of SBHBs. These objects became prime targets for further follow-up observations. After obtaining new observations over several years using ground-based telescopes, we identify J095036.75512838.1 (hereafter J0950) as a promising SBHB candidate.
1.2 Possible Interpretations of a Shifted Broad Emission Line and Scope of this Paper
Here we explore several interpretations for the apparent broad emission line shift of J0950. These include: (a) a binary supermassive black hole interpretation, which is the hypothesis originally considered in previous works on this object (Section 1.1), (b) a dust cloud outside the BLR obscuring parts of it as it moves across the line of sight, thereby producing asymmetric broad emission lines that mimic radial velocity shifts; (c) outflows, where the delays in the arrival times of ejecta moving towards and away from the observer result in apparent blueshifts and redshifts of the broad emission lines across different epochs; (d) a recoiling BH that has been kicked out of, and subsequently falls back into, the gravitational potential well of its host galaxy; and (e) a perturbation in a disk-like BLR that can produce a broad emission line peak that shifts in velocity as the perturbation precesses. We revisit these interpretations for further consideration in the Section 5.2.
The goal of this paper is to present the available data and a thorough study of the variability of the broad H line, and to explore the SBHB interpretation in detail. It is the first in a series of papers that will investigate possible scenarios for the behavior of this object through observational tests and/or modeling. In Section 2, we present the properties of the target and available spectroscopic and photometric data. Section 3 describes the methods for the velocity measurements, simulations of the observed profile change of the broad H emission line, and the relative velocities resulting from the H shift measurements. In Section 4, we present the analysis of the light curves. We assess the different interpretations of the observed behavior in J0950’s spectra and provide a detailed discussion of the SBHB interpretation in Section 5. Section 6 summarizes our results and notes future work. We assume a standard cosmology with , , and (bennett_cosmology2014) throughout this work.111We calculated luminosity distance using Ned Wright’s cosmology calculator (wright2006cosmology), available at https://www.astro.ucla.edu/~wright/CosmoCalc.html
| Telescope | Aperture | Observation | S/Nc |
| Instrumenta | Date (UT) | ||
| SloanSDSS | d=3′′ fiber | 2002.05.15 | 36 |
| HETLRS | slit | 2010.05.08 | 42 |
| HETLRS | slit | 2011.04.08 | 33 |
| SloanBOSS | d=2′′ fiber | 2015.01.20 | 33 |
| PalDBSP | slit | 2019.04.09 | 30 |
| HETLRS2 | 2020.02.02 | 19 | |
| HETLRS2 | 2020.04.02 | 19 | |
| HETLRS2 | 2021.04.05 | 36 | |
| KeckLRIS | slit | 2021.04.13 | 51 |
| HETLRS2 | 2022.03.25 | 36 | |
| HETLRS2 | 2022.04.03 | 21 | |
| HETLRS2 | 2022.12.31 | 28 | |
| HETLRS2 | 2024.03.30 | 28 |
Note. — a SDSS = Sloan Digital Sky Survey spectrograph, LRS = Low Resolution Spectrograph, BOSS = Baryon Oscillation Spectroscopic Survey spectrograph, DBSP = Double Spectrograph.
b The effective aperture size used in spectral extraction. LRS2 adjusts the fiber contributions based on an aperture set to 2.5 times the seeing value.
c The signal-to-noise ratio (S/N), estimated in the continuum around 5460 Å.
2 Properties of the Target and Available Data
2.1 Properties of the Target
J0950 is one of 88 SBHB candidates selected by eracleous2012large from a sample of 15,900 Sloan Digital Sky Survey (SDSS) quasars. The candidates were identified via velocity offsets of their broad H emission lines from the redshift defined by their [O iii] 4959,5007 narrow lines. Its redshift, measured from the rest wavelength of [O iii] 5007, is , which implies a luminosity distance of 1067.9 Mpc. Its absolute and apparent V-band magnitudes are 21.75 and 18.35 respectively. The short-term flux variability properties of the continuum and broad H emission of J0950 are much like those observed in normal quasars in the same redshift and luminosity range (runnoe2015large), making it appear to be a normal quasar apart from its offset broad-line.
2.2 Available Spectra
The current spectroscopic dataset for J0950 consists of spectra taken with the SDSS 2.5m telescope, the Palomar Hale 5m Telescope (Pal), the Keck 10m telescope, and the Hobby-Eberly 11m Telescope (HET) on the dates shown in Table 1 and Figure 1. The calibration process for the early spectra is described in Section 5.1 of eracleous2012large and summarized in Section 2.1 of runnoe2015large. The recent spectra from the HET were taken with the second-generation Low-Resolution Spectrograph (LRS2; lrs2_2016SPIE.9908E..4CC), which is fed by a fiber bundle. The LRS2 spectra were reduced and calibrated with the Panacea222https://github.com/grzeimann/Panacea package, written by Greg Zeimann. The reduction consists of bias and dark subtraction, fiber tracing and extraction as well as wavelength calibration, fiber-to-fiber normalization, source detection and extraction, and flux calibration. Starting with those spectra, we carried out the following corrections:
The observed flux density of each spectrum was divided by the continuous atmospheric extinction curve to recover the spectrum above the Earth’s atmosphere.
The flux density was corrected for Galactic interstellar extinction using the schlegel1998maps dust maps333The dust map was accessed using sfdmap (codebase: https://github.com/kbarbary/sfdmap) and extinction (barbary_k_2016) python packages. recalibrated by schlafly2011measuring, and the extinction law of fitzpatrick1999correcting.
The wavelength scale of each spectrum was converted from air to vacuum using the formula provided by key, which is an inverse transformation of the vacuum-to-air conversion given by morton2000atomic.
To verify the absolute wavelength calibration (hence the velocity stability) of the spectra, we used the [O iii] 4959,5007 doublet, which originates on large scales and serves as an internal wavelength standard. We found that, relative to the 2002 SDSS spectrum the subsequent spectra had shifts between and km s-1, as determined using the cross-correlation method described in Section 3. We corrected these shifts to adjust the spectral alignment. The residual uncertainty after the alignment process corresponds to km s-1, which is set by how well we can measure the shift in the [O iii] doublet via cross-correlation.
2.3 Available Light Curves
J0950 has archival optical and infrared (IR) light curves from
the Catalina Real-Time Transient Survey (CRTS; drake2009crts), the Near-Earth Object Wide-Field Infrared Survey Explorer (NEOWISE; wright2010wise; mainzer2014wise), and the Zwicky Transient Facility (ZTF; masci2019ztf; bellm2019ztf; graham2019ztf). Figure 2 shows the light curves from these facilities and in their respective filters.
We obtained the CRTS light curves (unfiltered visible light observations) from the second data release of The Catalina Surveys.444CRTS data access: http://nesssi.cacr.caltech.edu/DataRelease/CSDR2.html
Both NEOWISE (W1 and W2 filters) light curves and ZTF (g-, r-, and i-band) light curves from the 23rd data release were retrieved from the NASA/IPAC Infrared Science Archive (IRSA) archive.555NEOWISE data access: https://doi.org/10.26131/IRSA144
ZTF data access: https://doi.org/10.26131/IRSA598
The light curves span approximately 19 years in total. However, the segments covered by any single instrument are typically shorter than 10 years.
3 Analysis of Spectra
We aimed to measure velocity shifts of the broad H line in J0950 across different epochs. To do this, we used two main approaches: (a) cross correlating the H profiles directly in the observed spectra and (b) fitting the spectra with parametric models and cross correlating those in the models. In both cases, we also isolated the broad H line after decomposing the spectra to measure its velocity shifts independently of other spectral features. This required performing a spectral decomposition to remove contaminating components as needed. Our procedure yielded four distinct sets of relative velocity shifts of the broad H lines.
In Section 3.1, we describe the spectral decomposition and provide central moments and line properties measured from the model broad H lines. Section 3.2 outlines the cross-correlation technique and the resulting velocity shift measurements in both the data and models. In Sections 3.3 and 3.4, we perform simulations to investigate the effects of the observed variability of the broad H profile on the shift measurements. Section 3.5 summarizes the different sets of relative velocity shifts and explains which measurements we consider reliable and adopt for the subsequent radial velocity analysis. Throughout the above analyses we describe how we carry out experiments to determine uncertainties and, in Section 3.6 we summarize those uncertainties.
3.1 Spectral Decomposition and Measurements of the Broad H Line Profiles
We used the PyQSOfit quasar spectrum fitting code (guo2018pyqsofit) to model the spectra and decompose the host galaxy starlight spectrum, quasar continuum, Fe ii lines, broad and narrow He ii 4687 and H 4863 emission lines, and the [O iii] 4959,5007 doublet. The He ii feature is very weak in all the spectra and, although included in the fits, contributes negligibly to the overall models. To perform the decomposition, we restricted the spectral range to 4400–5500 Å and used only the least-squares routine for fitting. Each narrow emission line was modeled with a single Gaussian component, with the centroids and widths of the narrow lines tied together, reflecting their common narrow-line region origin. The broad H line was modeled with four Gaussian components to capture any profile asymmetries and to better fit the less well-defined profiles. The [O iii] wing components were modeled with a Gaussian each and tied to each other in both velocity and width. The width of the Fe ii emission was allowed to vary between 1200–18,000 km s-1. An example decomposition of the 2010 HET spectrum is shown in Figure 3.
The resulting models are shown in Appendix A, Figure 14. From these models, we isolated the model broad H profiles, as shown in Figure 15 (left panel). Starting from the central moment of a line profile, , where and are the individual wavelength and flux densities, and is a normalization constant, we calculated several line-shape metrics. The centroid was computed as . From this, we determined the skewness coefficient, , and the kurtosis coefficient, . We report these quantities along with the FWHM and integrated flux in Table 2. To quantify the uncertainties, we selected the model of the J0950 spectrum with median S/N (2015 HET spectrum; see Table 1), and generated 30 realizations by adding Gaussian noise with a constant standard deviation set by the measured S/N. For each realization, we modeled the broad H line and measured its central moments and line properties. We report the standard deviation of these measurements across the realizations as representative uncertainties for the central moments and line properties of the broad H line in all spectra of J0950.
The broad H profile recovered from a decomposition can be sensitive to the Fe ii fit. To assess this effect, we varied the width of the Fe ii lines in the decomposition process and examined its impact on the modeled broad H profile. We selected a subset of spectra that represent the range of observed broad H profiles of J0950. For each spectrum, we fixed the Fe ii width to be narrower or broader by up to 5000 km s-1 from the best-fit value, ensuring that each fixed width provided a reasonable fit while allowing all other parameters to vary. In each case, we omitted the model Fe ii and all other spectral features from the full model spectrum to isolate the broad H profile. We describe the implications of this Fe ii modeling experiment in Section 3.2.2, where we present the methodology for shift measurements.
| Observation | Centroid | Skewness | Kurtosis | FWHM | Integrated |
| Date (UT) | (km s-1)a | Coefficient | Coefficient | (km s-1) | Flux b |
| 2002.05.15 | 7.50 | ||||
| 2010.05.08 | 10.10 | ||||
| 2011.04.08 | 9.70 | ||||
| 2015.01.20 | 8.80 | ||||
| 2019.04.09 | 11.80 | ||||
| 2020.02.02 | 9.70 | ||||
| 2020.04.02 | 9.00 | ||||
| 2021.04.05 | 9.50 | ||||
| 2022.03.25 | 9.10 | ||||
| 2022.04.03 | 11.80 | ||||
| 2022.12.31 | 9.10 | ||||
| 2024.03.30 | 9.00 |
Note. — a Centroid shifts are expressed relative to the rest wavelength of H.
b Integrated fluxes are in units of erg s-1 cm-2.
3.2 Measuring Broad H Velocity Shifts
The technique used throughout this work to measure shifts of the broad H emission lines of J0950 is the cross-correlation method described in eracleous2012large (see also shen2013constraining). In summary, the method involves first taking two spectra, and linearly scaling one spectrum to match the other in its broad-line flux and continuum. The linear scaling minimizes the difference between the spectra so that the velocity offsets of their broad lines may be compared faithfully. To measure the relative broad H shift, one spectrum is held fixed while the other is progressively shifted in wavelength, and the at each shift is calculated in a window containing only the H profile. The shift value corresponding to the minimum is then adopted. Figure 9 in eracleous2012large demonstrates applications of the cross-correlation method.
In this work, each broad H velocity shift represents the average of two relative shift measurements between broad H profiles: one with the 2002 spectrum profile held constant while the profile of a later spectrum was shifted for cross-correlation, and another with the later spectrum held constant. Below we describe how we cross correlated the broad H lines in both the data and corresponding models to quantify sources of random and systematic errors.
3.2.1 Broad H Velocity Shifts in the Data
We measured velocity shifts of the broad H using two different approaches and compared the results. First, we cross correlated the full spectra in Figure 4, following the same procedure in eracleous2012large, after modeling and subtracting the host galaxy contribution with PyQSOfit (guo2018pyqsofit). In the second approach, we cross correlated the isolated broad H emission lines (Figure 5), obtained by following the spectral decomposition method adopted by other works (shen2013constraining; shen2019sloan; guo2019constraining). To isolate the broad H profiles, we first fit the spectra of J0950 as described in Section 3.1. After fitting, all components except the broad H were subtracted from each spectrum in Figure 4 to get the isolated broad H profiles shown in Figure 5.
The goal of measuring shifts using the full spectra and the isolated broad H emission lines is to check if the two methods yield consistent results. Figure 6 shows the relative velocities measured from the full spectra (red dotted line) and the isolated H lines (red circles), along with shaded regions and error bars representing 68% confidence intervals. We excluded the shift measurement from the 2021 Keck spectrum because a piece of the continuum near the broad H, which is essential for reliable spectral decomposition and cross-correlation, is missing due to the dichroic cutoff. Nonetheless, excluding this spectrum has a negligible effect as we have a measurement from an HET spectrum taken at approximately the same time.
As a consistency check, we applied the measured H shifts to four of our spectra that also include broad H (see Figure 1) and verified by visual inspection that those shifts bring the broad H profile into proper alignment. We do not use H as a kinematic tracer because it cannot be observed with H in a single exposure with our primary telescope, the HET, and because H at the redshift of J0950 is contaminated by strong sky lines that ultimately reduce the S/N and undermine our ability to use the narrow lines in its vicinity as a velocity standard. As a second consistency check, we cross-correlated successive spectra and verified that we obtained a shift equivalent to the difference of shifts to the first spectrum. Interestingly, among the observed spectra obtained within two months or less of each other in 2020 and 2022 (see Figure 4), there is a velocity difference of 80 km s-1 between the two spectra in 2022. Such random velocity variations on short timescales are often caused by subtle changes in the broad H line profile, which can be difficult to discern by eye. These velocity fluctuations are manifestations of radial velocity “jitter”, which we discuss further in Sections 3.6 and 5.3.
While the relative velocity shifts from the two methods increase uniformly, the shifts from cross-correlating the isolated H are systematically higher than those from the full spectra at higher velocities. This difference is primarily an artifact caused by the presence of Fe ii emission lines in the full spectra, as discussed in Section 3.2.2.
3.2.2 Broad H Velocity Shifts from Parametric Models of the Data
Using the spectral decomposition method described in Section 3.1, we generated models for the full spectra of J0950. From those, we isolated the broad H models. Models of the full spectra without the contribution from the host galaxy, and isolated broad H lines are shown in Appendix A. Following the methodology in Section 3.2.1, we measured shifts of the broad H using both the full spectra and the isolated broad H emission lines, this time analyzing the models instead of the data.
We first added synthetic Gaussian noise to each spectral model, as demonstrated in Appendix A and then measured the relative shifts of the broad H emission lines by cross correlating the noisy models. To quantify the measurement uncertainty, we generated 100 realizations of the noise for each model spectrum and repeated the cross-correlation. We took the median and 68% confidence interval of the resulting distribution as the final shift measurement and corresponding uncertainty, respectively.
Cross-correlating the broad H lines with the two approaches––using models of the full spectra and using models of the broad H profiles only––yielded two sets of relative velocity shifts, shown as a black dotted line and black circles in Figure 6. The two sets of measurements are qualitatively the same, with the velocity shift increasing monotonically in both the full and isolated broad H models. Nevertheless, we performed tests to determine which set of measurements is more reliable. Specifically, we cross correlated each isolated broad H line of the data in Figure 5 with its corresponding model (see Appendix A), which, as expected, yielded zero shift. However, including the Fe ii lines such that the cross-correlation is between an isolated broad H line from the data and its model of broad H plus Fe ii lines, gave non-zero shift values. Therefore, we conclude that the presence of the Fe ii complex introduces a systematic shift in the cross-correlation results. Moreover, the relative velocity shifts derived from the models (black dotted line and black circles in Figure 6) allow us to estimate the measurement uncertainty by carrying out multiple realizations of the noise. For these reasons, we adopt the relative velocity measurements based on the model broad H lines throughout this work. We also perform the same analysis on the relative velocity measurements from the isolated broad H lines of the data, and compare the results (see Section 3.5).
Furthermore, we used the Fe ii width experiment described in Section 3.1 to test how the broadening of the Fe ii template affects the relative velocity shifts obtained by cross correlating the isolated broad H lines. The aim was to assess how uncertainties in the Fe ii modeling would impact shift measurements. We measured the relative velocity shifts under two scenarios described in Section 3.1: one where the Fe ii was modeled as narrower, and another where it was modeled as broader than the best-fit width. The resulting shifts differ by about 10–75 km s-1, with the larger differences corresponding to the broader H profiles seen in the later spectra of J0950. However, these shifts are all consistent within the measurement uncertainties. The effect of varying the width of the Fe ii lines is small because the peak of the broad H, which is well captured in the modeling, contributes more strongly to cross-correlation measurements than the surrounding, less prominent features. Additionally, the Fe ii model fits the region under the broad H used for cross-correlation much better than it fits the strong Fe ii multiplets on either side (see the residuals in Figure 3).
3.3 Simulating the Profile Change of the Broad H
The profile of the broad H line of J0950 is observed to change such that it becomes broader over time. There is also a 20% stochastic variability of the integrated line flux––a small change that is not obvious from a visual inspection of the spectra in Figure 5; hidden by the fact that the broad H width is increasing.
Profile variability is common in normal quasars, but has been known to produce large uncertainties or even invalidate shift measurements (e.g., shen2013constraining; liu2014constraining; runnoe2017large). To test the effects of the observed profile variability of J0950 on our relative shift measurements, we simulated a spectral series based on the model of the broad H line of the 2002 SDSS spectrum. First, we convolved the model with Gaussian kernels of increasing width, and applied velocity shifts to the resulting profiles. Gaussian kernels offer a simple approximation without built-in physical assumptions, and allow us to emulate the evolution of the line profile. As we demonstrate in the next paragraph, using Gaussian kernels does not affect our shift measurements. Following the procedure in Section 3.2.2, we injected synthetic Gaussian noise into each simulated profile.
We show the simulated spectra in Figure 7, with the broadest spectrum having a full width at half maximum (FWHM) comparable to that of the last 2022 spectrum. We then carried out the cross-correlation procedure described in Section 3.2.2 to find the shift of each simulated profile relative to the 2002 broad H profile. The goal of this exercise was to see if cross-correlation can recover the shifts inserted into the simulated profiles. The results are shown in Figure 8, where each shift value corresponds to the median and 68% confidence interval of the distribution resulting from carrying out multiple realizations of noise before cross-correlation. Figure 8 confirms that the input shifts are reliably recovered despite the broadening of the profiles. From this simulation we also obtain an estimate of the typical uncertainty in the velocity shifts arising from the increasing width of the H profile with time: 80 km s-1. It is important to note that these simulated profiles are more round-topped than the actual observed profiles, all of which appear to have cusps (see Figure 5). Therefore, if cross-correlation can recover the shifts of profiles that are more round-topped than those of the data, then shift measurements made for the data are also reliable.
3.4 Simulating Breathing of the Broad-Line Region
We also explored whether the observed variation of J0950’s broad H line can be explained by “breathing” of the BLR (e.g., peterson2011masses; wang2020sloan). Breathing is the effect in which changes in the accretion-powered continuum of a quasar affect the extent of the zone in the BLR that can emit a line efficiently, hence the widths of the resulting broad lines become broader as those lines become fainter. To examine if the broad H behavior of J0950 can be replicated by breathing, and if that has an impact on our shift measurements, we repeated the exercise of Section 3.3 with the difference that we simulated the spectra by applying breathing effects. Specifically, we used a relationship derived under the premise that the observed continuum flux and width of the broad H line are connected to the mass of the BH by (bentz2009radius; peterson2011masses). Since is constant, we can write
where and are the integrated flux and width of the broad H emission respectively. and are the flux and width of the profile resulting from the convolution of the broad H line with a wider Gaussian kernel. Figure 9 shows the sequence of spectra that simulate symmetric profile variability expected from BLR breathing. The shift values recovered from cross-correlation are shown in Figure 10.
This exercise leads us to conclude the following: (a) the decline in integrated flux with increasing width of the broad lines in this simulation (Figure 9) is more drastic than that observed in the broad H line of J0950 (compare Figure 5). Moreover, as noted at the beginning of Section 3.3, the broad H flux of this object exhibits only a 20% stochastic variability––a behavior inconsistent with breathing since the FWHM increases systematically by a factor of over the monitoring period. (b) It is clear from Figure 10 that known shifts from the simulation can be measured despite the effects of breathing, which leads us to infer that symmetric profile variability, in general, has a negligible effect on the shift measurements.
3.5 The Radial Velocity Curve
In Sections 3.2.13.2.2 we obtained four different sets of relative velocity shifts by cross-correlating the broad H lines in: (a) the full spectra of the data, (b) the isolated broad H lines of the data, (c) the full spectra models, and (d) the isolated broad H models. In Section 3.2.2, we established that the presence of Fe ii lines in the full spectra (both data and models), introduces a systematic shift in the cross-correlation results, making shift measurements from the full spectra less reliable.
In view of the above, we proceeded only with the relative shift measurements based on the isolated broad H lines (Figure 5), and their corresponding models (Appendix A). We added to these shift values the peak offset measurement of the broad H of the 2002 spectrum. To measure the peak offset (following a method similar to eracleous2012large), we defined a fixed region around the peak and repeatedly fitted it with a Gaussian, each time varying the wavelength range used for fitting within that region. We took the median peak velocity as the best estimate of the true value and the 68% confidence interval of the resulting distribution of peak values as the measurement uncertainty. It is important to note that this method of locating the peak is sensitive to asymmetries within the broad emission line. Therefore, the Gaussian fitting region must be restricted to be near the peak, excluding significant asymmetries.
The addition of the peak offset from the 2002 spectrum provides the absolute peak offsets of the broad lines, as listed in Table 3. The measurements agree within 3, with average differences around 2. We return to these measurements in Section 5.3 to discuss the implications.
| Observation | Data H | Model H | ||
| Date (UT) | (km s-1) a | 95% b | (km s-1) c | 95% b |
| 2002.05.15 | () | () | ||
| 2010.05.08 | () | () | ||
| 2011.04.08 | () | () | ||
| 2015.01.20 | () | () | ||
| 2019.04.09 | () | () | ||
| 2020.02.02 | () | () | ||
| 2020.04.02 | () | () | ||
| 2021.04.05 | () | () | ||
| 2022.03.25 | () | () | ||
| 2022.04.03 | () | () | ||
| 2022.12.31 | () | () | ||
| 2024.03.30 | () | () |
Note. — Section 3.5 describes how the absolute shifts were obtained. In addition to the tabulated uncertainties, there is a systematic uncertainty of km s-1 from Fe ii subtraction (Section 3.2.2), a systematic uncertainty of km s-1 due to the increase in the width of the broad H profile with time (Section 3.3), and a 200 km s-1 jitter (Section 5.3).
a Absolute shifts and 68% confidence intervals of the isolated, broad H lines of the data shown in Figure 5.
b 95% confidence levels for the absolute shifts.
c Absolute shifts and 68% confidence intervals of the isolated, model broad H shown in Appendix A.
d H peak offset of the 2002 SDSS spectrum.
3.6 A Summary of Uncertainties
Throughout Section 3, we measured velocity shifts of the broad H line and did a series of tests and simulations to characterize the uncertainties in those measurements. The formal uncertainties from cross correlation are described in Section 3.2 and are listed in Table 3; they are between 40 and 60 km s-1. A systematic uncertainty of about 60 km s-1 arises from the modeling and subtraction of the Fe ii complex (Section 3.2.2). From the simulation in Section 3.3, we estimated an uncertainty of 80 km s-1 due to the increasing width of the broad H line over time. To account for short-timescale radial velocity fluctuations, such as the example in Section 3.2.1), we adopted a jitter of 200 km s-1 (discussed further in Section 5.3) for use in the radial velocity analyses presented later. The jitter is the dominant source of uncertainty.
4 Analysis of Light Curves
We investigated the optical and infrared (IR) light curves of J0950 using all publicly available data from the Catalina Real-Time Transient Survey (CRTS), the Near-Earth Object Wide-Field Infrared Survey Explorer (NEOWISE), and the Zwicky Transient Facility (ZTF). Although the available spectroscopic data span a longer time period than these light curves, we nonetheless analyzed the full extent of the available photometric data, shown in Figure 2.
While there is no apparent periodicity in the light curves, we computed the Lomb-Scargle Periodogram666We utilized the LombScargle feature from the astropy.timeseries python package. for the data from each facility separately, as well as for the combined light curve, with each dataset normalized to the CRTS light curve by its median magnitude. The period search was limited to periods up to 60% of the total time span of each dataset. Periods longer than the dataset duration are undetectable, and those longer than half the time span are unreliable, as they would not allow for the observation of at least two full cycles. In the periodograms, we set significance levels corresponding to 0.1% and 1% false alarm probabilities (corresponding to 99.9% and 99% confidence respectively), ensuring that any meaningful signal stands significantly above both the noise floor and these thresholds. The only significant signals appear at short periods corresponding to the observational cadences, around 1 yr.
In addition to the periodograms, we modeled each light curve with a flat line corresponding to its 3-clipped mean magnitude, and computed residuals normalized by the measurement uncertainties, as shown in Figure 11. The distributions of residuals normalized by the error bars appear approximately symmetric, though with a broader spread than a standard normal distribution, likely due to intrinsic stochastic variability of the quasar (e.g., vaughan2016false; liu2018did). The fit of a flat line in Figure 11 and the lack of any any periodic signal in the periodograms, leads us to conclude that there is no evidence of periodicity in the light curves. The light curves of J0950 are revisited in Section 5.5.
5 Interpretation, Discussion, and Speculation
5.1 Summary of Main Results from the Data
The quasar J0950 was suggested as a spectroscopic SBHB candidate based on the large velocity shift of its broad H line. Subsequent observations of this object reveal a progressive and systematic shift of its broad H line over time. It is particularly noteworthy that the velocity of the line changed sign during its evolution. The profile shape of the broad H line also changed between observations.
We used a cross-correlation technique to measure the velocity shifts of the broad H lines, applying it to both the observed line profiles and corresponding models. In each case, we used the full spectra of J0950 as well as the isolated H emission lines obtained through spectral decomposition. The relative velocity shifts from all approaches show a monotonic increase from negative to positive velocities. After extensive testing, we adopted the relative velocities derived from cross-correlating models of the isolated broad H lines as the most reliable. We also carried out the same analysis on the isolated broad H lines of the data in order to quantify systematic uncertainties. The choice of the isolated broad H models mitigates the systematic errors introduced by Fe ii lines, and the velocity measurements from these models provide an estimate of the measurement uncertainties through multiple noise realizations.
In addition to displaying a velocity shift, the broad H line profile of J0950 gradually broadened over time. This increase in width does not compromise our ability to measure shifts but introduces an additional uncertainty comparable in magnitude to the cross-correlation measurement uncertainties. The observed broadening is not consistent with the “breathing” behavior of a broad-line region.
Finally, we analyzed available light curves from CRTS, NEOWISE (W1 and W2 filters), and ZTF (g, r, and i bands). The light curves display a low level of apparently stochastic variability with no evidence for periodicity.
5.2 Assessment of Different Interpretations for the Observed Velocity Shifts
In this section, we consider the interpretations listed in Section 1.2 for the observed shift of the broad H line of J0950. We discuss those interpretations and evaluate whether they are physically plausible and applicable, before contemplating whether they can be subjected to statistical tests.
The broad H velocity shifts in the spectra of J0950 have increased steadily over the decades it has been monitored (Section 3.5). Such a trend resembles closely the expected signature of a long-period (several decades) binary supermassive black hole with a BLR associated with one of the two objects. The shifts in the broad H are then associated with the changes of the projected bulk velocity of the BLR over a portion of an orbital cycle. Since the time scale and general pattern of the radial velocity variations is consistent with expectations (e.g., eracleous2012large; runnoe2017large), we explore this hypothesis further by fitting a Keplerian orbit model to the observed radial velocity curve in Section 5.3 and considering the implications in Sections 5.4 and 5.5.
As a dust cloud outside the BLR and in a Keplerian orbit crosses our line of sight, it can obscure one side of the BLR first, then block a significant portion of the innermost BLR before obscuring the other side. Once the cloud moves out of our line of sight to the BLR, the full region becomes visible again. In a spectroscopic time series, this would result in the alternating appearance of the blueshifted and redshifted sides of a broad emission line. When the cloud is blocking the innermost BLR, the observed broad emission line appears centered, narrow and has a small integrated flux, since only the low-velocity gas in the outer parts of the BLR is visible. When the cloud is out of our line of sight to the BLR, a much stronger, broader line profile is observed because the full velocity field is visible. Therefore, a moving dust cloud can produce asymmetric broad emission lines and create apparent velocity shifts by selectively obscuring parts of the velocity field. However, the evolution of J0950’s broad H line is not consistent with this scenario. While we do observe blueshifted and redshifted broad-line peaks in the data, the dust-obscuration scenario is unlikely, since the spectrum in which the broad H line appears centered and relatively symmetric (2015 SDSS spectrum; Figure 4) is not extra strong and broad, as would be expected if the dust cloud were out of our line of sight to the BLR, nor narrow and weak, as would be expected if dust were temporarily obscuring the innermost parts of the BLR. Moreover, the observed variations of the broad H flux are only 20%.
We can also estimate the crossing time for an obscuring object orbiting the BLR by following the method of lamassa2015. For a supermassive BH with a BLR radius of pc (see estimates in Section 5.4), it would take years for a dust cloud in Keplerian orbit at that radius to cross the BLR. Given that the broad H line of J0950 exhibits a significant blueshift and redshift in the span of years, this timescale discrepancy makes it difficult to reconcile the observations with the dust-obscuration scenario. Furthermore, matching the observed timescale of about 20 years would require a BH mass , which is physically implausible (e.g., thomas201617; inayoshi2016there; mcconnell2012dynamical).
Another possible source of shifted broad emission lines is outflows. In the case of bipolar outflows (e.g., zheng1990outflows; veilleux1991outflows; zheng1991outflows) where the ejection is symmetric and expands outwards, the observer would first see the blueshifted side, followed by the redshifted side after a time delay. In this scenario, both the red and blue sides of a broad emission line must be visible. The only way for the blue side to appear first, then disappear, and then the red side appear, as we observe in J0950, would be if the blue side were to turn off exactly as the red side turns on. This would require exquisite fine tuning. Even with such fine tuning, the bipolar outflow scenario cannot explain the symmetric 2015 line profile centered at .
The possibility that J0950 is a recoiling BH is unlikely on several grounds. A recoiling BH, after being ejected, could reach a maximum distance from the galactic center up to several kpc (e.g., blecha2016recoil) before falling back and oscillating around the nucleus. However, the timescale for one full oscillation cycle is at least years (komossa&meritt2008recoil), which is far longer than the duration over which J0950 has been observed. Furthermore, although recoil speeds in excess of several thousand are possible in principle (e.g., healy2023recoil), the probability of observing a kick velocity that is from a supermassive binary merger is negligibly small; most recoil velocities are on the order of a few hundred or less (dotti2010dual).777Recoils from three-body interactions, can reach 1000 km s-1 for the lightest supermassive BH in a three-body interaction but are typically on the order of a few hundred (hoffman2006recoil). The likelihood of detecting kick velocities much greater than the escape speed of the host galaxy is even lower. In the case of J0950, where we observe velocities as high as , that probability is negligibly small (blecha2016recoil). Most critically, the radial velocity curve of J0950 mostly shows consistent positive acceleration as well as a change in sign––a behavior that is inconsistent with the expected dynamics of a recoiling BH.
There are known objects that show a single displaced peak with a shoulder, similar to the earliest spectrum of J0950. Examples include Mrk 668 and 3C 227, both of which have broad emission lines with a displaced peak and shoulder (marziani1993peculiar; gezari2007long). They have been monitored, more sparsely than J0950, and some do show radial velocity variations, but none have the pattern we see in J0950 (especially not the change of sign of the radial velocity curve). Non-axisymmetric disk models (e.g., storchi2017double; schimoia2017evolution) have been used to try to explain their irregular, non-periodic radial velocity curves. These models entail a “perturbation”, a bright spot or spiral arm, orbiting in the inner region of a circular, disk-like BLR, or a precessing elliptical disk (gezari2007long; lewis2010long; storchi2017double). This scenario can cause a broad emission line with two peaks, whose strengths alternate like a seesaw as the perturbation moves and enhances one peak after another. The smoothly increasing radial velocity variation of J0950 differs from the variations of these objects and cannot be explained by the non-axisymmetric BLR models described above. Nonetheless, the broader framework of a perturbed BLR may provide a plausible explanation for J0950’s behavior. Since other objects within this framework have shown substantial changes in their radial velocity over time, continued monitoring of J0950 is essential to see whether its trend persists or evolves (see the discussion of 3C 390.3 in Section 4.2 of eracleous1997rejection). Further testing of the single perturbed BLR scenario for J0950 will require developing a detailed physical model, which is out of the scope for this paper.
J0950 is an unusual object. Among quasars, only double-peaked emitters show behavior even remotely resembling the large monotonic velocity variations seen in J0950. doan2020improved present the longest monitoring campaign for double-peaked emitters; only two objects (3C 332 and 3C 390.3) displayed fairly monotonic velocity changes over most of their monitoring periods. Since double-peaked emitters make up only 3% of all quasars (strateva2003doublepeaked), such large monotonic variations are extremely rare––roughly 0.4% of quasars (2/14 of extensively monitored double-peaked emitters 3%). Importantly, none of these objects show a change in sign in their radial velocity curves. Therefore, the velocity changes observed in J0950 cannot be directly attributed to phenomena seen in other quasars.
Following the discussion above, we now explore in more detail whether the binary supermassive black hole interpretation is consistent with the available data. As noted earlier we will examine the single, perturbed BLR interpretation in detail in the future. We consider the remaining scenarios implausible and do not discuss them further.
5.3 A Supermassive Binary Black Hole Scenario for the Radial Velocity Curve
The radial velocity curve of J0950 based on the fourth column of Table 3 is shown in Figure 12. It shows a steady increase of 3500 km s-1 over time; the most striking behavior is that the broad line region of this quasar has undergone radial velocity variations in a manner resembling Keplerian motion in a binary. There certainly are potential caveats, which we discuss in Section 5.2. But the radial velocity curve of J0950 so far is consistent with binary motion, which makes it a strong SBHB candidate.
We increased the measurement uncertainties of the radial velocity curve to capture the effects of “jitter” from regular quasar variability (guo2019constraining; doan2020improved). Quasars can have perturbations or asymmetries within the broad line region that manifest as small, stochastic radial velocity variations, or jitter, occurring over a timescale of the order of a year or less (e.g., barth2015lick). This subtle effect is directly measurable in two of J0950’s spectra (see Section 3.2.1). doan2020improved empirically determined that jitter can be characterized as fluctuations around a smooth radial velocity curve, distributed approximately according to a Gaussian distribution (see also guo2019constraining). The standard deviation of the distribution is on the scale of a few hundred km s-1. Following the methodology from that work, we added a jitter of 200 km s-1 in quadrature to the uncertainties in each radial velocity measurement. The jitter dominates over the measurement errors and is the primary source of uncertainty in the radial velocity at each epoch.
To fit the radial velocity curve shown in Figure 12, we used radvel (fulton2018radvel), a package for fitting Keplerian orbits to radial velocity time series. The best-fitting orbital parameters, determined from the radial velocity curve fit in the observed frame, are provided in Table 4. The free parameters in the fitting are the period (), the time of conjunction (), the natural log of the semi-amplitude (ln ), and and , from which the eccentricity () and argument of periapsis () are derived. Given the exploratory nature of fitting the velocity curve of J0950 as a binary and the absence of prior knowledge about the system, we adopted broad, physically motivated priors based on the radial velocity curve. Specifically, we assumed a near-circular orbit as a starting point and adopted Gaussian priors on the period and time of conjunction: a mean period of yr (ensuring that the 1- range does not fall below the 22 yr span of radial velocity measurements available), and a time of conjunction yr. For a circular orbit, the conjunction––the moment when the lower-mass, or secondary BH, passes along the line of sight between the observer and the primary BH––would be expected near that epoch. A hard bound of was applied to the semi-amplitude to restrict unphysical amplitudes, and an eccentricity prior of was imposed. The Markov chain Monte Carlo (MCMC) was run with 20 walkers for 2500 steps each, and the chains converged reliably after steps (Gelman-Rubin statistic 1.01) following an initial burn-in. The resulting posteriors were well-constrained and consistent with the observed signal (see Appendix B for a corner plot), indicating that the priors were non-restrictive.
Since the radial velocity baseline of 22 yr covers only a fraction of the 33-yr orbit, the period, time of conjunction and the semi-amplitude, become partially degenerate. Therefore, the posterior distribution at longer periods is strongly influenced by the assumed prior. For much longer periods, the radial velocity curve would look almost the same over the observed 22 yr window. Likewise, the semi-amplitude can shift slightly depending on the adopted jitter value. These effects are expected for long period systems.
The median values of the posterior distributions derived from the MCMC analysis are reported in Table 4. The uncertainties are 68% confidence limits. Table 4 also provides the orbital parameters from performing the same analysis on the radial velocity curve obtained from the absolute shifts of the isolated, broad H lines of the data (provided in the second column of Table 3). The results of the two different fits capture the magnitude of the systematic uncertainties. The uncertainties encompass a wide range of possible Keplerian orbit models, such as those illustrated by the light orange lines in Figure 12. Observing a turnover in the radial velocity curve within the next few years would significantly constrain the possible models.
| Parameter | Model H a | Data H a |
| Free Parameters of the Model | ||
| Period, (yr)b | ||
| Time of conjunction, | ||
| Natural Log of Semi-amplitude, | ||
| c | ||
| c | ||
| Derived Parameters | ||
| Eccentricity, | ||
| Tangent of argument of periapsis, | ||
| Velocity semi-amplitude, (km s-1) | ||
| Minimum Reduced d | 0.2 | 0.3 |
Note. —
a All uncertainties correspond to 68% confidence intervals.
b The period reported here is in the frame of the observer.
c Finding eccentricity and argument of periapsis by fitting and is computationally more efficient in radvel than fitting and directly.
d Reduced with 7 degrees of freedom; a small value due to error bars increased by jitter.
5.4 Exploration of the Supermassive Binary Black Hole Scenario and Discussion
A lower limit on the total mass of the SBHB in J0950 was initially estimated at M⊙ by runnoe2017large. We obtained limits on the mass assuming that the bolometric luminosity does not exceed the Eddington luminosity. The bolometric luminosity was estimated by taking the luminosity of J0950’s 2002 spectrum––the brightest state of this object––at 5100 Å, and applying a bolometric correction factor of 10.3 (richards2006spectral). This yields a mass of . We estimated another lower mass limit, taking maximum inclination () and a mass ratio of , in the following Keplerian velocity equation,
This equation was derived under the assumption of an unequal mass binary system, in which the less massive, or secondary, BH is active and its BLR emits the observed H line (see arguments in eracleous2012large; runnoe2015large). Using all combinations explored by the MCMC algorithm for the period , the velocity semi-amplitude , and eccentricity , the inferred 68%-confidence lower limit on total mass is and the 99% lower limit is . These inferred masses do not necessarily correspond to the lower limits on , , and . All the lower mass limits above are reasonable in the sense that they are less than , a threshold regarded as unphysical. This threshold comes about because the most massive BHs measured are of the order of and because BHs are not expected to accrete material rapidly enough to surpass this mass (e.g., thomas201617; inayoshi2016there; mcconnell2012dynamical). In line with this argument, the minimum possible inclination for J0950 is . Using the distributions of period and mass, applying Kepler’s law gives 68% and 99% lower limits on the orbital semi-major axis of 0.008 pc and 0.006 pc respectively.
If J0950 is a binary, we attribute the offset broad H emission lines to a BLR associated with the lower mass, or secondary, BH. In this scenario, the tidal interaction with the companion BH may limit the size of the BLR. To explore that possibility, we estimate the “characteristic radius” of the BLR888This is the length that corresponds to the observed time delay, between fluctuations of continuum and a broad line, . using the empirical relation from bentz2013 between the characteristic radius and AGN luminosity at 5100 Å,
with and . We applied this relation even though it is appropriate only for a single BH and does not account for the dynamical effects in a binary system to get a characteristic radius of the BLR of pc. This value is somewhat higher than our inferred lower limit on the binary semi-major axis of pc. The extent of the line-emitting region in the BLR is possibly 45 times greater than (see discussion in runnoe2015large). Since we only have a lower limit on the orbital semi-major axis, there appears to be no conflict between this separation and the characteristic radius of the BLR.
If J0950 does indeed harbor a binary, extrapolating from the fit to its radial velocity curve, it should start turning over soon. It is therefore critical that we observe this object in the near future, at a rate of at least one spectrum per year to see if this prediction is borne out. It is also important to note that our implementation of radial velocity jitter discussed in Section 3.5 is simple, and more sophisticated prescriptions of jitter should be developed and employed. The description of jitter would impact the uncertainties on the orbital parameters that are inferred from fitting a radial velocity curve. Regardless of whether or not J0950 is a binary, it still displays an exceptional behavior that warrants detailed study.
5.5 On the Detectability of Doppler Boosting in the Light Curves
The light curves of J0950 exhibit neither substantial nor periodic variability, as discussed in Section 4. In addition to the tests in that section, we phase-folded the light curves using the inferred 33 yr period of J0950, binned the magnitude values by phase, and calculated the median magnitudes and spread within each phase bin. We fit the phase-folded, binned light curves with a flat line corresponding to the mean of each light curve, and found reduced values 5, consistent with no strong periodic structure. Nonetheless, we sought to estimate whether Doppler boosting could produce variability signatures that are either large enough to eventually emerge in the light curves, or too subtle that they remain undetectable.
Considering a SBHB system, in which a BH carrying the BLR beams light as it orbits, the observed specific flux depends on the angle between the beam’s velocity vector and the observer’s line of sight, through the relativistic Doppler factor , where and . For a single source, the observed flux scales as , where is the spectral index (see, for example, dorazio2023review).
Beaming is most pronounced when the binary is in quadrature. In such a configuration the angle is determined only by the inclination of the binary orbit so that , depending on whether the active black hole is moving towards or away from the observer ( is the velocity semi-amplitude obtained from the fit to the radial velocity curve). We can therefore express the ratio of the maximum to minimum observed flux as,
Adopting km s-1 and from our spectral fits, . This corresponds to a magnitude difference of ––a very small variation, far below the level of variability seen in the light curves of this object. Thus, in the interpretation of J0950 as a binary, the expected Doppler boosting signatures are likely buried in noise and too weak to be confidently detected. Still, the observed variability does not contradict the binary model.
5.6 Speculation: Evolution of J0950 in the GW landscape
Given our findings above, it is interesting to speculate about the expected evolution of the GW signal from a system like J0950 in the characteristic strain versus frequency diagram. Figure 13 illustrates the path of such a system as its eccentricity evolves from 0.65 to approximately zero, assuming that GW emission is the dominant mechanism driving orbital evolution. This assumption does not account for the ongoing accretion in the system, which implies continued interaction with the surrounding gas. Gas dynamics could influence the timescale of orbital decay and may sustain a non-zero eccentricity for an extended period. Figure 13 shows three different harmonics of the GW signal. The steps involved in producing the components of the figure are outlined in Appendix C. Here, we predict the overall evolution of the system across the PTA and LISA bands using theoretical sensitivity curves by hazboun2019 and robson2019, respectively. For the purposes of this illustration we adopt representative values of the mass and mass ratio, as indicated in Figure 13.
Initially, the characteristic strain, , of each higher-order harmonic, and , is strong. However, as the orbit circularizes, the harmonic becomes dominant, while the other harmonics decay quickly, albeit at different rates. As the eccentricity approaches zero, the becomes the only significant harmonic. To evolve the system further, up to the merger, we take the point of coalescence as approximately the separation at the innermost stable circular orbit (ISCO; the last stable circular orbit beyond which the BHs will merge). For a non-spinning (Schwarzschild) BH, the ISCO separation is given by (where is the mas of the more massive, or primary, BH). The path of the system from circularization of the orbit to ISCO, for varying configurations of total mass and mass ratios, is indicated by arrows in Figure 13, with the tip of the arrows representing the merger.
Finally, we estimate the evolutionary time scales for the GW signal from the system, adopting a characteristic binary mass of 10. Using the binary decay lifetime from peters1964, it will take 55 Myr for this object to merge. The details of this calculation are given in Appendix C. The time the system will spend in the PTA frequency band spans from when the harmonic (initially the strongest at the system’s current eccentricity, ) first enters the band, to the point where the harmonic intersects the PTA sensitivity curve. The system will spend 54 Myr in this band. Furthermore, it will take a time comparable to its overall decay lifetime of 55 Myr for it to evolve from its current eccentricity to the low-frequency end of the LISA band. Once in the LISA band, it will merge within 1 month.
It is clear from the exercise above that, if J0950 is a binary, it might eventually be detectable by PTA experiments. When it merges at the end of its evolution, it may be detectable by experiments similar to LISA. Therefore, J0950 could serve as an example of the systems producing the stochastic background detected by PTAs. Moreover, if there are other low-redshift systems with properties similar to those adopted above that are now emitting in the PTA band, we will have a chance of detecting them in the near future.
6 Conclusions
The spectroscopic supermassive black hole binary candidate J09505128 (abbreviated J0950) was identified in 2012 by the large velocity shift of its broad H line. Continued spectroscopic monitoring spanning 22 years, using the SDSS, Palomar Hale, Keck, and Hobby-Eberly telescopes, has revealed a systematic 3500 km s-1 shift in the broad H emission line over time, including a reversal in the sign of the velocity. Using a cross-correlation technique, we measured the velocity shifts in both the observed spectra and their model reconstructions, including versions with the isolated broad H component after spectral decomposition. All measurement approaches indicate a monotonic increase from negative to positive velocities.
Through tests and simulations, we identified sources of uncertainty of the same magnitude as those from cross correlation, including those arising from the subtraction Fe ii during spectral decomposition, as well as from the increasing width of the broad H line between observations. The largest and dominant source of uncertainty is jitter, which we combined with our measurement uncertainties to account for the intrinsic stochastic variability of quasars. Alongside the spectroscopic analyses, we examined the available light curves from CRTS, NEOWISE, and ZTF, and found no evidence of periodicity.
We considered several interpretations for the observed changes in broad emission line of J0950 and disfavored BLR obscuration by a dust cloud, outflows, and a recoiling black hole. The possibility that a SBHB or a perturbed, disk-like BLR could produce the broad emission line shifts remain viable. The latter interpretation requires additional data and specialized modeling for further tests that we defer to future work.
To evaluate the SBHB interpretation further, we fit the observed radial velocity curve with an orbital model. The fit yields a period of 33 years (in the frame of the observer) and an eccentricity of 0.65. We estimate lower limits on the semi-major axis and black hole mass of approximately 10-2 pc and , respectively. Under the binary interpretation of J0950, we used the orbital parameter estimates to explore whether Doppler boosting signatures could be detected in the light curves, finding that such signatures would be too weak to observe. We also used the orbital parameters to explore the evolution of a hypothetical GW signal from J0950. If the system is indeed a binary, its current expected signal falls in the detection band of PTA experiments, while its eventual merger signal would fall within the LISA detection band. We conclude that the quasar J0950 is an exceptional object. Regardless of whether it is a supermassive binary candidate, it warrants extensive monitoring to better characterize its properties and determine its true nature.
Acknowledgements
We thank the anonymous referee for their detailed comments that helped us improve the analysis of the data and presentation of the results. N.N.M acknowledges the support of the Penn State Science Achievement Graduate Fellowship, and helpful discussions with Laura Duffy, Meghan Delamer, Lucas Brefka, and Evan Fitzmaurice at Penn State. M.E. and N.N.M acknowledge support from the National Science Foundation under grant AST-2205720. J.C.R. and C.D. acknowledge support from the National National Foundation (NSF) under grant AST-2205719. T.B. acknowledges support from the National Science Foundation under grant AST-2307278 and from the Research Corporation for Science Advancement under award CS-SEED-2023-008. J.S is supported by an NSF Astronomy and Astrophysics Postdoctoral Fellowship under award AST-2202388, and acknowledges previous support from the JPL RTD program. M.C. is funded by the European Union (ERC-StG-2023, MMMonsters, 101117624). The work of D.S. and J.L. was carried out at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with NASA. We thank S. George Djorgovski for allowing us to use some of his time allocation on Keck to obtain spectra of J0950 on 04/13/2021.
The Low Resolution Spectrograph 2 (LRS2) was developed and funded by the University of Texas at Austin McDonald Observatory and Department of Astronomy, and by Pennsylvania State University. We thank the Leibniz-Institut fur Astrophysik Potsdam (AIP) and the Institut fur Astrophysik Goettingen (IAG) for their contributions to the construction of the integral field units. We acknowledge the Texas Advanced Computing Center (TACC) at The University of Texas at Austin for providing high performance computing, visualization, and storage resources that have contributed to the results reported within this paper.
Funding for the SDSS has been provided by the Alfred P. Sloan Foundation, the Participating Institutions, the National Science Foundation, the U.S. Department of Energy, the National Aeronautics and Space Administration, the Japanese Monbukagakusho, the Max Planck Society, and the Higher Education Funding Council for England. The SDSS Web Site is http://www.sdss.org/.
The SDSS is managed by the Astrophysical Research Consortium for the Participating Institutions. The Participating Institutions are the American Museum of Natural History, Astrophysical Institute Potsdam, University of Basel, University of Cambridge, Case Western Reserve University, University of Chicago, Drexel University, Fermilab, the Institute for Advanced Study, the Japan Participation Group, Johns Hopkins University, the Joint Institute for Nuclear Astrophysics, the Kavli Institute for Particle Astrophysics and Cosmology, the Korean Scientist Group, the Chinese Academy of Sciences (LAMOST), Los Alamos National Laboratory, the Max-Planck-Institute for Astronomy (MPIA), the Max-Planck-Institute for Astrophysics (MPA), New Mexico State University, Ohio State University, University of Pittsburgh, University of Portsmouth, Princeton University, the United States Naval Observatory, and the University of Washington.
Funding for SDSS IV has been provided by the Alfred P. Sloan Foundation, the U.S. Department of Energy Office of Science, and the Participating Institutions. SDSS-IV acknowledges support and resources from the Center for High Performance Computing at the University of Utah. The SDSS website is www.sdss4.org.
SDSS-IV is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration including the Brazilian Participation Group, the Carnegie Institution for Science, Carnegie Mellon University, Center for Astrophysics | Harvard & Smithsonian, the Chilean Participation Group, the French Participation Group, Instituto de Astrofísica de Canarias, The Johns Hopkins University, Kavli Institute for the Physics and Mathematics of the Universe (IPMU) / University of Tokyo, the Korean Participation Group, Lawrence Berkeley National Laboratory, Leibniz Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Astrophysik (MPA Garching), Max-Planck-Institut für Extraterrestrische Physik (MPE), National Astronomical Observatories of China, New Mexico State University, New York University, University of Notre Dame, Observatário Nacional / MCTI, The Ohio State University, Pennsylvania State University, Shanghai Astronomical Observatory, United Kingdom Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Oxford, University of Portsmouth, University of Utah, University of Virginia, University of Washington, University of Wisconsin, Vanderbilt University, and Yale University.
The NANOGrav collaboration receives support from National Science Foundation (NSF) Physics Frontiers Center award numbers 1430284 and 2020265, the Gordon and Betty Moore Foundation, NSF AccelNet award number 2114721, an NSERC Discovery Grant, and CIFAR. Part of this research was carried out at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration.
The CSS survey is funded by the National Aeronautics and Space Administration under Grant No. NNG05GF22G issued through the Science Mission Directorate Near-Earth Objects Observations Program. The CRTS survey is supported by the U.S. National Science Foundation under grants AST-0909182.
This publication makes use of data products from the Near-Earth Object Wide-field Infrared Survey Explorer (NEOWISE), which is a joint project of the Jet Propulsion Laboratory/California Institute of Technology and the University of California, Los Angeles. NEOWISE is funded by the National Aeronautics and Space Administration.
Based on observations obtained with the Samuel Oschin 48-inch Telescope at the Palomar Observatory as part of the Zwicky Transient Facility project. ZTF is supported by the National Science Foundation under Grant No. AST-1440341 and a collaboration including Caltech, IPAC, the Weizmann Institute for Science, the Oskar Klein Center at Stockholm University, the University of Maryland, the University of Washington, Deutsches Elektronen-Synchrotron and Humboldt University, Los Alamos National Laboratories, the TANGO Consortium of Taiwan, the University of Wisconsin at Milwaukee, and Lawrence Berkeley National Laboratories. Operations are conducted by COO, IPAC, and UW.
This research has made use of the NASA/IPAC Infrared Science Archive, which is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology.
Appendix A Side-by-Side Comparison of Models With and Without Noise for the Full Spectra and Isolated Broad H Lines of J0950
In this appendix, we show spectral decomposition models for the spectra of J0950. Each model in Figure 14 includes the quasar continuum, Fe ii lines, broad and narrow He ii 4687 (a weak line in every spectrum of J0950) and H 4863 emission lines, and the [O iii] 4959,5007 doublet (see Figure 3 for an annotated illustration of these features), while Figure 15 shows models of the isolated broad H lines. Also presented in both figures are noisy versions of the models. These model spectra are referenced in Section 3.2, where we quantify sources of random and systematic errors by cross correlating the broad H emission lines in the full and isolated spectral models.
Appendix B MCMC posterior distributions of orbital parameters from radial velocity fits of J0950
In Section 5.3, we performed MCMC fits to two radial velocity curves of J0950: one based on the absolute shifts of the isolated broad H models, and another based on the absolute shifts measured from the isolated broad H lines of the data. The fitted orbital parameters are the period (), the time of conjunction (), the natural log of the velocity semi-amplitude (ln ), and and . Figure 16 shows the corner plot of the posterior distributions from fitting the model-based velocity curve, overlaid with those from fitting the data-based curve.
Appendix C Evolution of a system like J0950 in the PTA and LISA gravitational-wave landscape
The characteristic strain in the detector frame of an eccentric binary at the th harmonic, can be determined using a simplified expression from breivik2020,
Here, is the chirp mass, is luminosity distance, is the detector frame GW frequency, is an eccentricity enhancement factor, and is a function involving a series of Bessel functions of the first kind, that captures the contribution of each harmonic and eccentricity to the gravitational radiation power spectrum (see peters&mathews1963).
To produce the evolutionary tracks of a system like J0950 for the different harmonics shown in Figure 13, we considered a range of eccentricities from 0.65 to approximately zero. For a given , the corresponding orbital separation can be obtained using an expression from peters1964,
where is calculated from the instantaneous and , which we take to be pc (semi-major axis for a binary mass of 10 and orbital period of 33 yr) and . Then for each separation, the GW frequency is,
The equations above were applied iteratively over a range of eccentricities at a specific . Once the orbit becomes circular, the harmonic dominates. The system was then evolved from the frequency where the eccentricity is approximately zero to the frequency at merger, maintaining a fixed . Recall that for this exercise, we defined “merger” as the separation at ISCO in Section 5.6.
The timescales provided in Section 5.6 for such a system to merge and its GW signal to be detected by PTA experiments and LISA, were computed using the decay lifetime equation given by peters1964,
where . For example, we found that the time it will take for such a system to merge, assuming it is an equal-mass 10 binary, is 55 Myr. This was determined by starting with an initial eccentricity of and a final eccentricity corresponding to the separation at ISCO. In Figure 13, the ISCO is indicated by the tip of the arrow of the harmonic. To determine the value of , we first took the GW frequency at this tip and calculated the corresponding separation by rearranging Eq. C3. In this particular example, the ISCO separation can also be directly calculated using . Then, we obtained by numerically evaluating in Eq. C2. Finally, the time to evolve from to was calculated via numerical integration.

### v4

@@
 We consider several interpretations for the observed changes, including a binary supermassive black hole, dust-cloud obscuration, outflows, a recoiling black hole, and a single perturbed, disk-like broad-line region.
-We disfavor all but the binary and perturbed broad-line region scenarios.
+We deem the binary and perturbed broad-line region scenarios to be physically plausible.
 The binary interpretation is the only one for which we can immediately compare a physical model to the available data.
@@
 The fit yields a period of 33 years (observed frame) and an eccentricity of 0.65, with lower limits on the semi-major axis and black hole mass of pc and , respectively.
-Thus, J09505128 is a compelling supermassive binary candidate.
+Thus, J09505128 is a binary candidate deserving further study.
 The single, perturbed broad-line region interpretation remains viable but requires additional observations and modeling for further evaluation.
 Continued monitoring is, therefore, essential.
-1 Introduction
-1.1 Background and Motivation
-Galaxy mergers are a frequently observed phenomenon, and given that most massive galaxies have central supermassive black holes (BHs; kormendy1995inward), supermassive black hole binaries (SBHBs) are thought to be an inevitable consequence of the merger process.
-begelman1980massive described how a SBHB evolves in a post-galaxy merger.
+I Introduction
+I.1 Background and Motivation
+Galaxy mergers are a frequently observed phenomenon, and given that most massive galaxies have central supermassive black holes (BHs; Kormendy and Richstone 1995), supermassive black hole binaries (SBHBs) are thought to be an inevitable consequence of the merger process.
+Begelman et al.
+(1980) described how a SBHB evolves in a post-galaxy merger.
 The evolution begins with the stellar bulges and respective BHs of the two galaxies sinking toward the center of the merger remnant.
-This process is initially driven by dynamical friction (chandrasekhar1943dynamical), which acts as the dominant mechanism for angular momentum loss for 109 yr for typical masses and can bring the BH pair closer.
-At a separation of pc, the BHs form a gravitationally bound binary (e.g., bogdanovic2022electromagnetic).
+This process is initially driven by dynamical friction (Chandrasekhar 1943), which acts as the dominant mechanism for angular momentum loss for 109 yr for typical masses and can bring the BH pair closer.
+At a separation of pc, the BHs form a gravitationally bound binary (e.g., Bogdanović et al.
+2022).
 Eventually, the binary “hardens” as its orbital velocity becomes comparable to the velocity dispersion of the stars.
 At this point, stellar scattering takes over as the dominant mechanism that further reduces the binary separation.
-Stellar scattering is effective so long as the number of stars with the right orbits to interact with the binary can be replenished (milosavljevic2003long; yu2005evolution).
+Stellar scattering is effective so long as the number of stars with the right orbits to interact with the binary can be replenished (Milosavljević and Merritt 2003; Yu et al.
+2005).
 If this reservoir of stars is depleted, the binary separation may slow its evolution or stall at a separation of pc (known as the “last parsec problem”).
-However, several models––ranging from angular momentum loss through interactions with a circumbinary accretion disk or gaseous cloud (e.g., armitage2002accretion; escala2004role; dotti_lisa2006; cuadra2009massive) to dynamical models that consider non-spherical and anisotropic stellar distributions (e.g., merritt2004chaotic; berczik2006efficient; khan2012mergers; khan2013supermassive; holley2015galaxy)––now suggest that there are a number of sufficiently efficient mechanisms for continued orbital decay.
+However, several models––ranging from angular momentum loss through interactions with a circumbinary accretion disk or gaseous cloud (e.g., Armitage and Natarajan 2002; Escala et al.
+2004; Dotti et al.
+2006; Cuadra et al.
+2009) to dynamical models that consider non-spherical and anisotropic stellar distributions (e.g., Merritt and Poon 2004; Berczik et al.
+2006; Khan et al.
+2012, 2013; Holley-Bockelmann and Khan 2015)––now suggest that there are a number of sufficiently efficient mechanisms for continued orbital decay.
 Once the binary separation reaches 10-2 pc, angular momentum loss via gravitational waves (GWs) becomes an efficient mechanism, leading the two BHs to coalesce within a Hubble time.
 Unambiguous observational evidence for SBHBs has been confined to the early stages of their evolution.
-Galaxy mergers and widely separated (kilo-parsec and sub-kilo-parsec scale) BH pairs are observed (e.g., comerford2009inspiralling; comerford20091; liu2011dual; ellison2017discovery; hou2019active; foord2020second; goulding2019dual; shen2021dual; koss2023dual), but bound SBHBs have yet to be imaged and resolved directly.
-An exception is the radio-loud candidate CSO 0402+379, directly imaged through radio interferometry with a projected separation of 7 pc (rodriguez2006compact; bansal2017constraining).
-The recent evidence of a stochastic GW background by pulsar timing array (PTA) experiments is compatible with the low-frequency signal expected from a population of SBHBs at the last stages of their evolution (agazie2023nanograv; agazie2023binary; antoniadis2023second; reardon2023search; lee2023searching; miles2025meerkat), strongly suggesting that SBHBs must evolve to merger, emitting GWs in the process.
+Galaxy mergers and widely separated (kilo-parsec and sub-kilo-parsec scale) BH pairs are observed (e.g., Comerford et al.
+2009a, b; Liu et al.
+2011; Ellison et al.
+2017; Hou et al.
+2019; Foord et al.
+2020; Goulding et al.
+2019; Shen et al.
+2021; Koss et al.
+2023), but bound SBHBs have yet to be imaged and resolved directly.
+An exception is the radio-loud candidate CSO 0402+379, directly imaged through radio interferometry with a projected separation of 7 pc (Rodriguez et al.
+2006; Bansal et al.
+2017).
+The recent evidence of a stochastic GW background by pulsar timing array (PTA) experiments is compatible with the low-frequency signal expected from a population of SBHBs at the last stages of their evolution (Agazie et al.
+2023a, b; EPTA Collaboration et al.
+2023; Reardon et al.
+2023; Xu et al.
+2023; Miles et al.
+2025), strongly suggesting that SBHBs must evolve to merger, emitting GWs in the process.
 This is promising for directly resolving nanohertz GWs (corresponding to orbital periods of tens of years) from individual SBHBs at sub-parsec separations.
-The upcoming Laser Interferometer Space Antenna (LISA; amaro2023astrophysics; colpi2024) is also expected to detect SBHB mergers, at higher frequencies corresponding to somewhat lower BH masses and later stages of SBHB evolution than those detected by the PTAs.
-SBHBs have also been invoked to explain a number of galaxy properties, including mass deficits in the cores of elliptical galaxies (e.g., milosavljevic2002cores; merritt2006cores), helical or precessing jets (e.g., roos1993precessing; romero2000preccessing), and X-shaped radio sources (e.g., merritt2002xshaped; gopalkrishna2003xshaped).
+The upcoming Laser Interferometer Space Antenna (LISA; Amaro-Seoane et al.
+2023; Colpi et al.
+2024) is also expected to detect SBHB mergers, at higher frequencies corresponding to somewhat lower BH masses and later stages of SBHB evolution than those detected by the PTAs.
+SBHBs have also been invoked to explain a number of galaxy properties, including mass deficits in the cores of elliptical galaxies (e.g., Milosavljević et al.
+2002; Merritt 2006), helical or precessing jets (e.g., Roos et al.
+1993; Romero et al.
+2000), and X-shaped radio sources (e.g., Merritt and Ekers 2002; Gopal-Krishna et al.
+2003).
 Finding unambiguous SBHBs at small separations ( 0.11 pc) would greatly enhance our understanding of their evolution and their effects on their host galaxies.
@@
 The two most commonly used methods, which we describe further below, rely on periodic photometric variability or regular spectroscopic (radial velocity) variability of the broad emission lines.
-In addition to these, a few other methods have been proposed that are appropriate for SBHBs at separations of – 1 pc: peculiar profiles and relative intensities of broad emission lines (e.g., montuori11; montuori12), proper motion measurements via infrared interferometry (e.g.
-dexter20), periodic modulation of the polarization of scattered light (e.g.
-dotti22), kinematic signatures in reverberation mapping signals (e.g., wang18), or blue-red asymmetric reverberation of the broad emission lines (e.g.
-dotti23).
-The photometric method involves looking for periodic optical flux variability in the light curves of quasars (e.g., graham2015systematic; charisi2016population; liu2016systematic) under the assumption that either one or both BHs are accreting, and that the accretion rate is regulated on the binary orbital period (d2013accretion).
-Periodic oscillations can also occur due to relativistic Doppler boosting from an orbiting active BH (dorazio2023review).
-A prominent photometric SBHB candidate is OJ 287, whose long-term light curve displayed outbursts separated on a 12-year period (e.g., valtonen2008Nature; valtonen2012; komossa2023oj27_1; komossa2023oj287_2).
-The use of periodic photometric variations to identify binaries, especially when only a small number of apparent orbital cycles is observed, however, is subject to ambiguities (liu2018did; witt2022period; robnik2024period), since quasars exhibit stochastic variability that can appear periodic (vaughan2016false; liu2016systematic; barth_stern2018; elbadry2025arXiv).
-The spectroscopic method looks for the displaced peaks of broad emission lines relative to the host galaxy rest frame in quasar spectra (e.g., gaskell1983quasars; gaskell1996evidence; gaskell1996supermassive)––this is the approach used in this paper.
+In addition to these, a few other methods have been proposed that are appropriate for SBHBs at separations of – 1 pc: peculiar profiles and relative intensities of broad emission lines (e.g., Montuori et al., 2011, 2012), proper motion measurements via infrared interferometry (e.g.
+Dexter et al., 2020), periodic modulation of the polarization of scattered light (e.g.
+Dotti et al., 2022), kinematic signatures in reverberation mapping signals (e.g., Wang et al., 2018), or blue-red asymmetric reverberation of the broad emission lines (e.g.
+Dotti et al., 2023).
+The photometric method involves looking for periodic optical flux variability in the light curves of quasars (e.g., Graham et al.
+2015; Charisi et al.
+2016; Liu et al.
+2016b) under the assumption that either one or both BHs are accreting, and that the accretion rate is regulated on the binary orbital period (D’Orazio et al.
+2013).
+Periodic oscillations can also occur due to relativistic Doppler boosting from an orbiting active BH (D’Orazio and Charisi 2023).
+A prominent photometric SBHB candidate is OJ 287, whose long-term light curve displayed outbursts separated on a 12-year period (e.g., Valtonen et al.
+2008, 2012; Komossa et al.
+2023b, a).
+The use of periodic photometric variations to identify binaries, especially when only a small number of apparent orbital cycles is observed, however, is subject to ambiguities (Liu et al.
+2018; Witt et al.
+2022; Robnik et al.
+2024), since quasars exhibit stochastic variability that can appear periodic (Vaughan et al.
+2016; Liu et al.
+2016b; Barth and Stern 2018; El-Badry et al.
+2025).
+The spectroscopic method looks for the displaced peaks of broad emission lines relative to the host galaxy rest frame in quasar spectra (e.g., Gaskell 1983, 1996b, 1996a)––this is the approach used in this paper.
 A single-peaked broad emission line that is significantly offset is a possible signature of bulk motion of a single active BH and its surrounding gas in a binary.
-The possibility that the double-peaked emitters of some quasars can arise from systems with two active BHs has been tested and rejected based on the lack of radial velocity variations that are indicative of orbital motion (eracleous1997rejection; liu2016doublepeak; doan2020improved).
-Rather, double-peaked emission lines are best explained as originating in a disk-like broad-line region around a single supermassive black hole (see, for example, eracleous2003radioloud; eracleous2009review, for summary of arguments).
-While the single-peaked offset line method is viable for SBHB searches, it suffers from the ambiguity that offsets of broad emission lines can also be caused by other effects, such as outflows from the accretion disk (e.g., popovic2012super) or large, non-axisymmetric perturbations of the broad-line region (BLR; e.g., gezari2007long; lewis2010long; storchi2017double; schimoia2017evolution).
+The possibility that the double-peaked emitters of some quasars can arise from systems with two active BHs has been tested and rejected based on the lack of radial velocity variations that are indicative of orbital motion (Eracleous et al.
+1997; Liu et al.
+2016a; Doan et al.
+2020).
+Rather, double-peaked emission lines are best explained as originating in a disk-like broad-line region around a single supermassive black hole (see, for example, Eracleous and Halpern, 2003; Eracleous et al., 2009, for summary of arguments).
+While the single-peaked offset line method is viable for SBHB searches, it suffers from the ambiguity that offsets of broad emission lines can also be caused by other effects, such as outflows from the accretion disk (e.g., Popović 2012) or large, non-axisymmetric perturbations of the broad-line region (BLR; e.g., Gezari et al.
+2007; Lewis et al.
+2010; Storchi-Bergmann et al.
+2017; Schimoia et al.
+2017).
 It is therefore necessary to have multiple observations over extended periods of time to evaluate SBHB candidates identified by this method.
-In order to select SBHB candidates based on their displaced broad emission lines, one approach is to search for 1000 km s-1 velocity offsets of the broad H emission lines relative to the narrow emission lines in the spectra of quasars (gaskell1983quasars; tsalmantza2011systematic; eracleous2012large; decarli2013nature; liu2014constraining).
-Using this method, eracleous2012large selected 88 SBHB candidates which have since undergone extensive studies to characterize their nature such as flux variability, radial velocity changes, and implied population properties (runnoe2015large; runnoe2017large; pflueger2018likelihood; nguyen2020pulsar).
-runnoe2017large used observations through 2011 and found 29 candidates that displayed steady, and in some cases, statistically significant radial velocity variations.
+In order to select SBHB candidates based on their displaced broad emission lines, one approach is to search for 1000 km s-1 velocity offsets of the broad H emission lines relative to the narrow emission lines in the spectra of quasars (Gaskell 1983; Tsalmantza et al.
+2011; Eracleous et al.
+2012; Decarli et al.
+2013; Liu et al.
+2014).
+Using this method, Eracleous et al.
+(2012) selected 88 SBHB candidates which have since undergone extensive studies to characterize their nature such as flux variability, radial velocity changes, and implied population properties (Runnoe et al.
+2015, 2017; Pflueger et al.
+2018; Nguyen et al.
+2020).
+Runnoe et al.
+(2017) used observations through 2011 and found 29 candidates that displayed steady, and in some cases, statistically significant radial velocity variations.
 Among these, three objects displayed monotonic changes consistent with the expected behavior of SBHBs.
@@
 After obtaining new observations over several years using ground-based telescopes, we identify J095036.75512838.1 (hereafter J0950) as a promising SBHB candidate.
-1.2 Possible Interpretations of a Shifted Broad Emission Line and Scope of this Paper
+I.2 Possible Interpretations of a Shifted Broad Emission Line and Scope of this Paper
 Here we explore several interpretations for the apparent broad emission line shift of J0950.
-These include: (a) a binary supermassive black hole interpretation, which is the hypothesis originally considered in previous works on this object (Section 1.1), (b) a dust cloud outside the BLR obscuring parts of it as it moves across the line of sight, thereby producing asymmetric broad emission lines that mimic radial velocity shifts; (c) outflows, where the delays in the arrival times of ejecta moving towards and away from the observer result in apparent blueshifts and redshifts of the broad emission lines across different epochs; (d) a recoiling BH that has been kicked out of, and subsequently falls back into, the gravitational potential well of its host galaxy; and (e) a perturbation in a disk-like BLR that can produce a broad emission line peak that shifts in velocity as the perturbation precesses.
-We revisit these interpretations for further consideration in the Section 5.2.
+These include: (a) a binary supermassive black hole interpretation, which is the hypothesis originally considered in previous works on this object (Section I.1), (b) a dust cloud outside the BLR obscuring parts of it as it moves across the line of sight, thereby producing asymmetric broad emission lines that mimic radial velocity shifts; (c) outflows, where the delays in the arrival times of ejecta moving towards and away from the observer result in apparent blueshifts and redshifts of the broad emission lines across different epochs; (d) a recoiling BH that has been kicked out of, and subsequently falls back into, the gravitational potential well of its host galaxy; and (e) a perturbation in a disk-like BLR that can produce a broad emission line peak that shifts in velocity as the perturbation precesses.
+We revisit these interpretations for further consideration in the Section V.2.
 The goal of this paper is to present the available data and a thorough study of the variability of the broad H line, and to explore the SBHB interpretation in detail.
 It is the first in a series of papers that will investigate possible scenarios for the behavior of this object through observational tests and/or modeling.
-In Section 2, we present the properties of the target and available spectroscopic and photometric data.
-Section 3 describes the methods for the velocity measurements, simulations of the observed profile change of the broad H emission line, and the relative velocities resulting from the H shift measurements.
-In Section 4, we present the analysis of the light curves.
-We assess the different interpretations of the observed behavior in J0950’s spectra and provide a detailed discussion of the SBHB interpretation in Section 5.
-Section 6 summarizes our results and notes future work.
-We assume a standard cosmology with , , and (bennett_cosmology2014) throughout this work.111We calculated luminosity distance using Ned Wright’s cosmology calculator (wright2006cosmology), available at https://www.astro.ucla.edu/~wright/CosmoCalc.html
+In Section II, we present the properties of the target and available spectroscopic and photometric data.
+Section III describes the methods for the velocity measurements, simulations of the observed profile change of the broad H emission line, and the relative velocities resulting from the H shift measurements.
+In Section IV, we present the analysis of the light curves.
+We assess the different interpretations of the observed behavior in J0950’s spectra and provide a detailed discussion of the SBHB interpretation in Section V.
+Section VI summarizes our results and notes future work.
+We assume a standard cosmology with , , and (Bennett et al., 2014) throughout this work.111We calculated luminosity distance using Ned Wright’s cosmology calculator (Wright, 2006), available at https://www.astro.ucla.edu/~wright/CosmoCalc.html
 | Telescope | Aperture | Observation | S/Nc |
@@
 c The signal-to-noise ratio (S/N), estimated in the continuum around 5460 Å.
-2 Properties of the Target and Available Data
-2.1 Properties of the Target
-J0950 is one of 88 SBHB candidates selected by eracleous2012large from a sample of 15,900 Sloan Digital Sky Survey (SDSS) quasars.
+II Properties of the Target and Available Data
+II.1 Properties of the Target
+J0950 is one of 88 SBHB candidates selected by Eracleous et al.
+(2012) from a sample of 15,900 Sloan Digital Sky Survey (SDSS) quasars.
 The candidates were identified via velocity offsets of their broad H emission lines from the redshift defined by their [O iii] 4959,5007 narrow lines.
@@
 Its absolute and apparent V-band magnitudes are 21.75 and 18.35 respectively.
-The short-term flux variability properties of the continuum and broad H emission of J0950 are much like those observed in normal quasars in the same redshift and luminosity range (runnoe2015large), making it appear to be a normal quasar apart from its offset broad-line.
-2.2 Available Spectra
+The short-term flux variability properties of the continuum and broad H emission of J0950 are much like those observed in normal quasars in the same redshift and luminosity range (Runnoe et al.
+2015), making it appear to be a normal quasar apart from its offset broad-line.
+II.2 Available Spectra
 The current spectroscopic dataset for J0950 consists of spectra taken with the SDSS 2.5m telescope, the Palomar Hale 5m Telescope (Pal), the Keck 10m telescope, and the Hobby-Eberly 11m Telescope (HET) on the dates shown in Table 1 and Figure 1.
-The calibration process for the early spectra is described in Section 5.1 of eracleous2012large and summarized in Section 2.1 of runnoe2015large.
-The recent spectra from the HET were taken with the second-generation Low-Resolution Spectrograph (LRS2; lrs2_2016SPIE.9908E..4CC), which is fed by a fiber bundle.
+The calibration process for the early spectra is described in Section 5.1 of Eracleous et al.
+(2012) and summarized in Section 2.1 of Runnoe et al.
+(2015).
+The recent spectra from the HET were taken with the second-generation Low-Resolution Spectrograph (LRS2; Chonis et al.
+2016), which is fed by a fiber bundle.
 The LRS2 spectra were reduced and calibrated with the Panacea222https://github.com/grzeimann/Panacea package, written by Greg Zeimann.
@@
 The observed flux density of each spectrum was divided by the continuous atmospheric extinction curve to recover the spectrum above the Earth’s atmosphere.
-The flux density was corrected for Galactic interstellar extinction using the schlegel1998maps dust maps333The dust map was accessed using sfdmap (codebase: https://github.com/kbarbary/sfdmap) and extinction (barbary_k_2016) python packages.
-recalibrated by schlafly2011measuring, and the extinction law of fitzpatrick1999correcting.
-The wavelength scale of each spectrum was converted from air to vacuum using the formula provided by key, which is an inverse transformation of the vacuum-to-air conversion given by morton2000atomic.
+The flux density was corrected for Galactic interstellar extinction using the Schlegel et al.
+(1998) dust maps333The dust map was accessed using sfdmap (codebase: https://github.com/kbarbary/sfdmap) and extinction (Barbary 2016) python packages.
+recalibrated by Schlafly and Finkbeiner (2011), and the extinction law of Fitzpatrick (1999).
 To verify the absolute wavelength calibration (hence the velocity stability) of the spectra, we used the [O iii] 4959,5007 doublet, which originates on large scales and serves as an internal wavelength standard.
-We found that, relative to the 2002 SDSS spectrum the subsequent spectra had shifts between and km s-1, as determined using the cross-correlation method described in Section 3.
+We found that, relative to the 2002 SDSS spectrum the subsequent spectra had shifts between and km s-1, as determined using the cross-correlation method described in Section III.
 We corrected these shifts to adjust the spectral alignment.
 The residual uncertainty after the alignment process corresponds to km s-1, which is set by how well we can measure the shift in the [O iii] doublet via cross-correlation.
-2.3 Available Light Curves
+II.3 Available Light Curves
 J0950 has archival optical and infrared (IR) light curves from
-the Catalina Real-Time Transient Survey (CRTS; drake2009crts), the Near-Earth Object Wide-Field Infrared Survey Explorer (NEOWISE; wright2010wise; mainzer2014wise), and the Zwicky Transient Facility (ZTF; masci2019ztf; bellm2019ztf; graham2019ztf).
+the Catalina Real-Time Transient Survey (CRTS; Drake et al.
+2009), the Near-Earth Object Wide-Field Infrared Survey Explorer (NEOWISE; Wright et al.
+2010; Mainzer et al.
+2014), and the Zwicky Transient Facility (ZTF; Masci et al.
+2019; Bellm et al.
+2019; Graham et al.
+2019).
 Figure 2 shows the light curves from these facilities and in their respective filters.
@@
 The light curves span approximately 19 years in total. However, the segments covered by any single instrument are typically shorter than 10 years.
-3 Analysis of Spectra
+III Analysis of Spectra
 We aimed to measure velocity shifts of the broad H line in J0950 across different epochs.
@@
 Our procedure yielded four distinct sets of relative velocity shifts of the broad H lines.
-In Section 3.1, we describe the spectral decomposition and provide central moments and line properties measured from the model broad H lines.
-Section 3.2 outlines the cross-correlation technique and the resulting velocity shift measurements in both the data and models.
-In Sections 3.3 and 3.4, we perform simulations to investigate the effects of the observed variability of the broad H profile on the shift measurements.
-Section 3.5 summarizes the different sets of relative velocity shifts and explains which measurements we consider reliable and adopt for the subsequent radial velocity analysis.
-Throughout the above analyses we describe how we carry out experiments to determine uncertainties and, in Section 3.6 we summarize those uncertainties.
-3.1 Spectral Decomposition and Measurements of the Broad H Line Profiles
-We used the PyQSOfit quasar spectrum fitting code (guo2018pyqsofit) to model the spectra and decompose the host galaxy starlight spectrum, quasar continuum, Fe ii lines, broad and narrow He ii 4687 and H 4863 emission lines, and the [O iii] 4959,5007 doublet.
+In Section III.1, we describe the spectral decomposition and provide central moments and line properties measured from the model broad H lines.
+Section III.2 outlines the cross-correlation technique and the resulting velocity shift measurements in both the data and models.
+In Sections III.3 and III.4, we perform simulations to investigate the effects of the observed variability of the broad H profile on the shift measurements.
+Section III.5 summarizes the different sets of relative velocity shifts and explains which measurements we consider reliable and adopt for the subsequent radial velocity analysis.
+Throughout the above analyses we describe how we carry out experiments to determine uncertainties and, in Section III.6 we summarize those uncertainties.
+III.1 Spectral Decomposition and Measurements of the Broad H Line Profiles
+We used the PyQSOfit quasar spectrum fitting code (Guo et al.
+2018) to model the spectra and decompose the host galaxy starlight spectrum, quasar continuum, Fe ii lines, broad and narrow He ii 4687 and H 4863 emission lines, and the [O iii] 4959,5007 doublet.
 The He ii feature is very weak in all the spectra and, although included in the fits, contributes negligibly to the overall models.
@@
 In each case, we omitted the model Fe ii and all other spectral features from the full model spectrum to isolate the broad H profile.
-We describe the implications of this Fe ii modeling experiment in Section 3.2.2, where we present the methodology for shift measurements.
+We describe the implications of this Fe ii modeling experiment in Section III.2.2, where we present the methodology for shift measurements.
 | Observation | Centroid | Skewness | Kurtosis | FWHM | Integrated |
@@
 b Integrated fluxes are in units of erg s-1 cm-2.
-3.2 Measuring Broad H Velocity Shifts
-The technique used throughout this work to measure shifts of the broad H emission lines of J0950 is the cross-correlation method described in eracleous2012large (see also shen2013constraining).
+III.2 Measuring Broad H Velocity Shifts
+The technique used throughout this work to measure shifts of the broad H emission lines of J0950 is the cross-correlation method described in Eracleous et al.
+(2012) (see also Shen et al., 2013).
 In summary, the method involves first taking two spectra, and linearly scaling one spectrum to match the other in its broad-line flux and continuum.
@@
 The shift value corresponding to the minimum is then adopted.
-Figure 9 in eracleous2012large demonstrates applications of the cross-correlation method.
+Figure 9 in Eracleous et al.
+(2012) demonstrates applications of the cross-correlation method.
 In this work, each broad H velocity shift represents the average of two relative shift measurements between broad H profiles: one with the 2002 spectrum profile held constant while the profile of a later spectrum was shifted for cross-correlation, and another with the later spectrum held constant.
 Below we describe how we cross correlated the broad H lines in both the data and corresponding models to quantify sources of random and systematic errors.
-3.2.1 Broad H Velocity Shifts in the Data
+III.2.1 Broad H Velocity Shifts in the Data
 We measured velocity shifts of the broad H using two different approaches and compared the results.
-First, we cross correlated the full spectra in Figure 4, following the same procedure in eracleous2012large, after modeling and subtracting the host galaxy contribution with PyQSOfit (guo2018pyqsofit).
-In the second approach, we cross correlated the isolated broad H emission lines (Figure 5), obtained by following the spectral decomposition method adopted by other works (shen2013constraining; shen2019sloan; guo2019constraining).
-To isolate the broad H profiles, we first fit the spectra of J0950 as described in Section 3.1.
+First, we cross correlated the full spectra in Figure 4, following the same procedure in Eracleous et al.
+(2012), after modeling and subtracting the host galaxy contribution with PyQSOfit (Guo et al.
+2018).
+In the second approach, we cross correlated the isolated broad H emission lines (Figure 5), obtained by following the spectral decomposition method adopted by other works (Shen et al.
+2013, 2019; Guo et al.
+2019).
+To isolate the broad H profiles, we first fit the spectra of J0950 as described in Section III.1.
 After fitting, all components except the broad H were subtracted from each spectrum in Figure 4 to get the isolated broad H profiles shown in Figure 5.
@@
 Such random velocity variations on short timescales are often caused by subtle changes in the broad H line profile, which can be difficult to discern by eye.
-These velocity fluctuations are manifestations of radial velocity “jitter”, which we discuss further in Sections 3.6 and 5.3.
+These velocity fluctuations are manifestations of radial velocity “jitter”, which we discuss further in Sections III.6 and V.3.
 While the relative velocity shifts from the two methods increase uniformly, the shifts from cross-correlating the isolated H are systematically higher than those from the full spectra at higher velocities.
-This difference is primarily an artifact caused by the presence of Fe ii emission lines in the full spectra, as discussed in Section 3.2.2.
-3.2.2 Broad H Velocity Shifts from Parametric Models of the Data
-Using the spectral decomposition method described in Section 3.1, we generated models for the full spectra of J0950.
+This difference is primarily an artifact caused by the presence of Fe ii emission lines in the full spectra, as discussed in Section III.2.2.
+III.2.2 Broad H Velocity Shifts from Parametric Models of the Data
+Using the spectral decomposition method described in Section III.1, we generated models for the full spectra of J0950.
 From those, we isolated the broad H models.
 Models of the full spectra without the contribution from the host galaxy, and isolated broad H lines are shown in Appendix A.
-Following the methodology in Section 3.2.1, we measured shifts of the broad H using both the full spectra and the isolated broad H emission lines, this time analyzing the models instead of the data.
+Following the methodology in Section III.2.1, we measured shifts of the broad H using both the full spectra and the isolated broad H emission lines, this time analyzing the models instead of the data.
 We first added synthetic Gaussian noise to each spectral model, as demonstrated in Appendix A and then measured the relative shifts of the broad H emission lines by cross correlating the noisy models.
@@
 For these reasons, we adopt the relative velocity measurements based on the model broad H lines throughout this work.
-We also perform the same analysis on the relative velocity measurements from the isolated broad H lines of the data, and compare the results (see Section 3.5).
-Furthermore, we used the Fe ii width experiment described in Section 3.1 to test how the broadening of the Fe ii template affects the relative velocity shifts obtained by cross correlating the isolated broad H lines.
+We also perform the same analysis on the relative velocity measurements from the isolated broad H lines of the data, and compare the results (see Section III.5).
+Furthermore, we used the Fe ii width experiment described in Section III.1 to test how the broadening of the Fe ii template affects the relative velocity shifts obtained by cross correlating the isolated broad H lines.
 The aim was to assess how uncertainties in the Fe ii modeling would impact shift measurements.
-We measured the relative velocity shifts under two scenarios described in Section 3.1: one where the Fe ii was modeled as narrower, and another where it was modeled as broader than the best-fit width.
+We measured the relative velocity shifts under two scenarios described in Section III.1: one where the Fe ii was modeled as narrower, and another where it was modeled as broader than the best-fit width.
 The resulting shifts differ by about 10–75 km s-1, with the larger differences corresponding to the broader H profiles seen in the later spectra of J0950.
@@
 Additionally, the Fe ii model fits the region under the broad H used for cross-correlation much better than it fits the strong Fe ii multiplets on either side (see the residuals in Figure 3).
-3.3 Simulating the Profile Change of the Broad H
+III.3 Simulating the Profile Change of the Broad H
 The profile of the broad H line of J0950 is observed to change such that it becomes broader over time.
 There is also a 20% stochastic variability of the integrated line flux––a small change that is not obvious from a visual inspection of the spectra in Figure 5; hidden by the fact that the broad H width is increasing.
-Profile variability is common in normal quasars, but has been known to produce large uncertainties or even invalidate shift measurements (e.g., shen2013constraining; liu2014constraining; runnoe2017large).
+Profile variability is common in normal quasars, but has been known to produce large uncertainties or even invalidate shift measurements (e.g., Shen et al.
+2013; Liu et al.
+2014; Runnoe et al.
+2017).
 To test the effects of the observed profile variability of J0950 on our relative shift measurements, we simulated a spectral series based on the model of the broad H line of the 2002 SDSS spectrum.
@@
 As we demonstrate in the next paragraph, using Gaussian kernels does not affect our shift measurements.
-Following the procedure in Section 3.2.2, we injected synthetic Gaussian noise into each simulated profile.
+Following the procedure in Section III.2.2, we injected synthetic Gaussian noise into each simulated profile.
 We show the simulated spectra in Figure 7, with the broadest spectrum having a full width at half maximum (FWHM) comparable to that of the last 2022 spectrum.
-We then carried out the cross-correlation procedure described in Section 3.2.2 to find the shift of each simulated profile relative to the 2002 broad H profile.
+We then carried out the cross-correlation procedure described in Section III.2.2 to find the shift of each simulated profile relative to the 2002 broad H profile.
 The goal of this exercise was to see if cross-correlation can recover the shifts inserted into the simulated profiles.
@@
 Therefore, if cross-correlation can recover the shifts of profiles that are more round-topped than those of the data, then shift measurements made for the data are also reliable.
-3.4 Simulating Breathing of the Broad-Line Region
-We also explored whether the observed variation of J0950’s broad H line can be explained by “breathing” of the BLR (e.g., peterson2011masses; wang2020sloan).
+III.4 Simulating Breathing of the Broad-Line Region
+We also explored whether the observed variation of J0950’s broad H line can be explained by “breathing” of the BLR (e.g., Peterson 2011; Wang et al.
+2020).
 Breathing is the effect in which changes in the accretion-powered continuum of a quasar affect the extent of the zone in the BLR that can emit a line efficiently, hence the widths of the resulting broad lines become broader as those lines become fainter.
-To examine if the broad H behavior of J0950 can be replicated by breathing, and if that has an impact on our shift measurements, we repeated the exercise of Section 3.3 with the difference that we simulated the spectra by applying breathing effects.
-Specifically, we used a relationship derived under the premise that the observed continuum flux and width of the broad H line are connected to the mass of the BH by (bentz2009radius; peterson2011masses).
+To examine if the broad H behavior of J0950 can be replicated by breathing, and if that has an impact on our shift measurements, we repeated the exercise of Section III.3 with the difference that we simulated the spectra by applying breathing effects.
+Specifically, we used a relationship derived under the premise that the observed continuum flux and width of the broad H line are connected to the mass of the BH by (Bentz et al.
+2009; Peterson 2011).
 Since is constant, we can write
@@
 This exercise leads us to conclude the following: (a) the decline in integrated flux with increasing width of the broad lines in this simulation (Figure 9) is more drastic than that observed in the broad H line of J0950 (compare Figure 5).
-Moreover, as noted at the beginning of Section 3.3, the broad H flux of this object exhibits only a 20% stochastic variability––a behavior inconsistent with breathing since the FWHM increases systematically by a factor of over the monitoring period.
+Moreover, as noted at the beginning of Section III.3, the broad H flux of this object exhibits only a 20% stochastic variability––a behavior inconsistent with breathing since the FWHM increases systematically by a factor of over the monitoring period.
 (b) It is clear from Figure 10 that known shifts from the simulation can be measured despite the effects of breathing, which leads us to infer that symmetric profile variability, in general, has a negligible effect on the shift measurements.
-3.5 The Radial Velocity Curve
-In Sections 3.2.13.2.2 we obtained four different sets of relative velocity shifts by cross-correlating the broad H lines in: (a) the full spectra of the data, (b) the isolated broad H lines of the data, (c) the full spectra models, and (d) the isolated broad H models.
-In Section 3.2.2, we established that the presence of Fe ii lines in the full spectra (both data and models), introduces a systematic shift in the cross-correlation results, making shift measurements from the full spectra less reliable.
+III.5 The Radial Velocity Curve
+In Sections III.2.1III.2.2 we obtained four different sets of relative velocity shifts by cross-correlating the broad H lines in: (a) the full spectra of the data, (b) the isolated broad H lines of the data, (c) the full spectra models, and (d) the isolated broad H models.
+In Section III.2.2, we established that the presence of Fe ii lines in the full spectra (both data and models), introduces a systematic shift in the cross-correlation results, making shift measurements from the full spectra less reliable.
 In view of the above, we proceeded only with the relative shift measurements based on the isolated broad H lines (Figure 5), and their corresponding models (Appendix A).
 We added to these shift values the peak offset measurement of the broad H of the 2002 spectrum.
-To measure the peak offset (following a method similar to eracleous2012large), we defined a fixed region around the peak and repeatedly fitted it with a Gaussian, each time varying the wavelength range used for fitting within that region.
+To measure the peak offset (following a method similar to Eracleous et al.
+2012), we defined a fixed region around the peak and repeatedly fitted it with a Gaussian, each time varying the wavelength range used for fitting within that region.
 We took the median peak velocity as the best estimate of the true value and the 68% confidence interval of the resulting distribution of peak values as the measurement uncertainty.
@@
 The measurements agree within 3, with average differences around 2.
-We return to these measurements in Section 5.3 to discuss the implications.
+We return to these measurements in Section V.3 to discuss the implications.
 | Observation | Data H | Model H | ||
@@
 Note.
-— Section 3.5 describes how the absolute shifts were obtained.
-In addition to the tabulated uncertainties, there is a systematic uncertainty of km s-1 from Fe ii subtraction (Section 3.2.2), a systematic uncertainty of km s-1 due to the increase in the width of the broad H profile with time (Section 3.3), and a 200 km s-1 jitter (Section 5.3).
+— Section III.5 describes how the absolute shifts were obtained.
+In addition to the tabulated uncertainties, there is a systematic uncertainty of km s-1 from Fe ii subtraction (Section III.2.2), a systematic uncertainty of km s-1 due to the increase in the width of the broad H profile with time (Section III.3), and a 200 km s-1 jitter (Section V.3).
 a Absolute shifts and 68% confidence intervals of the isolated, broad H lines of the data shown in Figure 5.
@@
 d H peak offset of the 2002 SDSS spectrum.
-3.6 A Summary of Uncertainties
-Throughout Section 3, we measured velocity shifts of the broad H line and did a series of tests and simulations to characterize the uncertainties in those measurements.
-The formal uncertainties from cross correlation are described in Section 3.2 and are listed in Table 3; they are between 40 and 60 km s-1.
-A systematic uncertainty of about 60 km s-1 arises from the modeling and subtraction of the Fe ii complex (Section 3.2.2).
-From the simulation in Section 3.3, we estimated an uncertainty of 80 km s-1 due to the increasing width of the broad H line over time.
-To account for short-timescale radial velocity fluctuations, such as the example in Section 3.2.1), we adopted a jitter of 200 km s-1 (discussed further in Section 5.3) for use in the radial velocity analyses presented later.
+III.6 A Summary of Uncertainties
+Throughout Section III, we measured velocity shifts of the broad H line and did a series of tests and simulations to characterize the uncertainties in those measurements.
+The formal uncertainties from cross correlation are described in Section III.2 and are listed in Table 3; they are between 40 and 60 km s-1.
+A systematic uncertainty of about 60 km s-1 arises from the modeling and subtraction of the Fe ii complex (Section III.2.2).
+From the simulation in Section III.3, we estimated an uncertainty of 80 km s-1 due to the increasing width of the broad H line over time.
+To account for short-timescale radial velocity fluctuations, such as the example in Section III.2.1), we adopted a jitter of 200 km s-1 (discussed further in Section V.3) for use in the radial velocity analyses presented later.
 The jitter is the dominant source of uncertainty.
-4 Analysis of Light Curves
+IV Analysis of Light Curves
 We investigated the optical and infrared (IR) light curves of J0950 using all publicly available data from the Catalina Real-Time Transient Survey (CRTS), the Near-Earth Object Wide-Field Infrared Survey Explorer (NEOWISE), and the Zwicky Transient Facility (ZTF).
@@
 In addition to the periodograms, we modeled each light curve with a flat line corresponding to its 3-clipped mean magnitude, and computed residuals normalized by the measurement uncertainties, as shown in Figure 11.
-The distributions of residuals normalized by the error bars appear approximately symmetric, though with a broader spread than a standard normal distribution, likely due to intrinsic stochastic variability of the quasar (e.g., vaughan2016false; liu2018did).
+The distributions of residuals normalized by the error bars appear approximately symmetric, though with a broader spread than a standard normal distribution, likely due to intrinsic stochastic variability of the quasar (e.g., Vaughan et al., 2016; Liu et al., 2018).
 The fit of a flat line in Figure 11 and the lack of any any periodic signal in the periodograms, leads us to conclude that there is no evidence of periodicity in the light curves.
-The light curves of J0950 are revisited in Section 5.5.
-5 Interpretation, Discussion, and Speculation
-5.1 Summary of Main Results from the Data
+The light curves of J0950 are revisited in Section V.5.
+V Interpretation, Discussion, and Speculation
+V.1 Summary of Main Results from the Data
 The quasar J0950 was suggested as a spectroscopic SBHB candidate based on the large velocity shift of its broad H line.
@@
 The light curves display a low level of apparently stochastic variability with no evidence for periodicity.
-5.2 Assessment of Different Interpretations for the Observed Velocity Shifts
-In this section, we consider the interpretations listed in Section 1.2 for the observed shift of the broad H line of J0950.
+V.2 Assessment of Different Interpretations for the Observed Velocity Shifts
+In this section, we consider the interpretations listed in Section I.2 for the observed shift of the broad H line of J0950.
 We discuss those interpretations and evaluate whether they are physically plausible and applicable, before contemplating whether they can be subjected to statistical tests.
-The broad H velocity shifts in the spectra of J0950 have increased steadily over the decades it has been monitored (Section 3.5).
+The broad H velocity shifts in the spectra of J0950 have increased steadily over the decades it has been monitored (Section III.5).
 Such a trend resembles closely the expected signature of a long-period (several decades) binary supermassive black hole with a BLR associated with one of the two objects.
 The shifts in the broad H are then associated with the changes of the projected bulk velocity of the BLR over a portion of an orbital cycle.
-Since the time scale and general pattern of the radial velocity variations is consistent with expectations (e.g., eracleous2012large; runnoe2017large), we explore this hypothesis further by fitting a Keplerian orbit model to the observed radial velocity curve in Section 5.3 and considering the implications in Sections 5.4 and 5.5.
+Since the time scale and general pattern of the radial velocity variations is consistent with expectations (e.g., Eracleous et al., 2012; Runnoe et al., 2017), we explore this hypothesis further by fitting a Keplerian orbit model to the observed radial velocity curve in Section V.3 and considering the implications in Sections V.4 and V.5.
 As a dust cloud outside the BLR and in a Keplerian orbit crosses our line of sight, it can obscure one side of the BLR first, then block a significant portion of the innermost BLR before obscuring the other side.
@@
 Moreover, the observed variations of the broad H flux are only 20%.
-We can also estimate the crossing time for an obscuring object orbiting the BLR by following the method of lamassa2015.
-For a supermassive BH with a BLR radius of pc (see estimates in Section 5.4), it would take years for a dust cloud in Keplerian orbit at that radius to cross the BLR.
+We can also estimate the crossing time for an obscuring object orbiting the BLR by following the method of LaMassa et al.
+(2015).
+For a supermassive BH with a BLR radius of pc (see estimates in Section V.4), it would take years for a dust cloud in Keplerian orbit at that radius to cross the BLR.
 Given that the broad H line of J0950 exhibits a significant blueshift and redshift in the span of years, this timescale discrepancy makes it difficult to reconcile the observations with the dust-obscuration scenario.
-Furthermore, matching the observed timescale of about 20 years would require a BH mass , which is physically implausible (e.g., thomas201617; inayoshi2016there; mcconnell2012dynamical).
+Furthermore, matching the observed timescale of about 20 years would require a BH mass , which is physically implausible (e.g., Thomas et al., 2016; Inayoshi and Haiman, 2016; McConnell et al., 2012).
 Another possible source of shifted broad emission lines is outflows.
-In the case of bipolar outflows (e.g., zheng1990outflows; veilleux1991outflows; zheng1991outflows) where the ejection is symmetric and expands outwards, the observer would first see the blueshifted side, followed by the redshifted side after a time delay.
+In the case of bipolar outflows (e.g., Zheng et al., 1990; Veilleux and Zheng, 1991; Zheng et al., 1991) where the ejection is symmetric and expands outwards, the observer would first see the blueshifted side, followed by the redshifted side after a time delay.
 In this scenario, both the red and blue sides of a broad emission line must be visible.
@@
 The possibility that J0950 is a recoiling BH is unlikely on several grounds.
-A recoiling BH, after being ejected, could reach a maximum distance from the galactic center up to several kpc (e.g., blecha2016recoil) before falling back and oscillating around the nucleus.
-However, the timescale for one full oscillation cycle is at least years (komossa&meritt2008recoil), which is far longer than the duration over which J0950 has been observed.
-Furthermore, although recoil speeds in excess of several thousand are possible in principle (e.g., healy2023recoil), the probability of observing a kick velocity that is from a supermassive binary merger is negligibly small; most recoil velocities are on the order of a few hundred or less (dotti2010dual).777Recoils from three-body interactions, can reach 1000 km s-1 for the lightest supermassive BH in a three-body interaction but are typically on the order of a few hundred (hoffman2006recoil).
+A recoiling BH, after being ejected, could reach a maximum distance from the galactic center up to several kpc (e.g., Blecha et al., 2016) before falling back and oscillating around the nucleus.
+However, the timescale for one full oscillation cycle is at least years (Komossa and Merritt, 2008), which is far longer than the duration over which J0950 has been observed.
+Furthermore, although recoil speeds in excess of several thousand are possible in principle (e.g., Healy and Lousto, 2023), the probability of observing a kick velocity that is from a supermassive binary merger is negligibly small; most recoil velocities are on the order of a few hundred or less (Dotti et al., 2010).777Recoils from three-body interactions, can reach 1000 km s-1 for the lightest supermassive BH in a three-body interaction but are typically on the order of a few hundred (Hoffman and Loeb, 2006).
 The likelihood of detecting kick velocities much greater than the escape speed of the host galaxy is even lower.
-In the case of J0950, where we observe velocities as high as , that probability is negligibly small (blecha2016recoil).
+In the case of J0950, where we observe velocities as high as , that probability is negligibly small (Blecha et al., 2016).
 Most critically, the radial velocity curve of J0950 mostly shows consistent positive acceleration as well as a change in sign––a behavior that is inconsistent with the expected dynamics of a recoiling BH.
 There are known objects that show a single displaced peak with a shoulder, similar to the earliest spectrum of J0950.
-Examples include Mrk 668 and 3C 227, both of which have broad emission lines with a displaced peak and shoulder (marziani1993peculiar; gezari2007long).
+Examples include Mrk 668 and 3C 227, both of which have broad emission lines with a displaced peak and shoulder (Marziani et al.
+1993; Gezari et al.
+2007).
 They have been monitored, more sparsely than J0950, and some do show radial velocity variations, but none have the pattern we see in J0950 (especially not the change of sign of the radial velocity curve).
-Non-axisymmetric disk models (e.g., storchi2017double; schimoia2017evolution) have been used to try to explain their irregular, non-periodic radial velocity curves.
-These models entail a “perturbation”, a bright spot or spiral arm, orbiting in the inner region of a circular, disk-like BLR, or a precessing elliptical disk (gezari2007long; lewis2010long; storchi2017double).
+Non-axisymmetric disk models (e.g., Storchi-Bergmann et al.
+2017; Schimoia et al.
+2017) have been used to try to explain their irregular, non-periodic radial velocity curves.
+These models entail a “perturbation”, a bright spot or spiral arm, orbiting in the inner region of a circular, disk-like BLR, or a precessing elliptical disk (Gezari et al., 2007; Lewis et al., 2010; Storchi-Bergmann et al., 2017).
 This scenario can cause a broad emission line with two peaks, whose strengths alternate like a seesaw as the perturbation moves and enhances one peak after another.
@@
 Nonetheless, the broader framework of a perturbed BLR may provide a plausible explanation for J0950’s behavior.
-Since other objects within this framework have shown substantial changes in their radial velocity over time, continued monitoring of J0950 is essential to see whether its trend persists or evolves (see the discussion of 3C 390.3 in Section 4.2 of eracleous1997rejection).
+Since other objects within this framework have shown substantial changes in their radial velocity over time, continued monitoring of J0950 is essential to see whether its trend persists or evolves (see the discussion of 3C 390.3 in Section 4.2 of Eracleous et al.
+1997).
 Further testing of the single perturbed BLR scenario for J0950 will require developing a detailed physical model, which is out of the scope for this paper.
@@
 Among quasars, only double-peaked emitters show behavior even remotely resembling the large monotonic velocity variations seen in J0950.
-doan2020improved present the longest monitoring campaign for double-peaked emitters; only two objects (3C 332 and 3C 390.3) displayed fairly monotonic velocity changes over most of their monitoring periods.
-Since double-peaked emitters make up only 3% of all quasars (strateva2003doublepeaked), such large monotonic variations are extremely rare––roughly 0.4% of quasars (2/14 of extensively monitored double-peaked emitters 3%).
+Doan et al.
+(2020) present the longest monitoring campaign for double-peaked emitters; only two objects (3C 332 and 3C 390.3) displayed fairly monotonic velocity changes over most of their monitoring periods.
+Since double-peaked emitters make up only 3% of all quasars (Strateva et al., 2003), such large monotonic variations are extremely rare––roughly 0.4% of quasars (2/14 of extensively monitored double-peaked emitters 3%).
 Importantly, none of these objects show a change in sign in their radial velocity curves.
@@
 We consider the remaining scenarios implausible and do not discuss them further.
-5.3 A Supermassive Binary Black Hole Scenario for the Radial Velocity Curve
+V.3 A Supermassive Binary Black Hole Scenario for the Radial Velocity Curve
 The radial velocity curve of J0950 based on the fourth column of Table 3 is shown in Figure 12.
 It shows a steady increase of 3500 km s-1 over time; the most striking behavior is that the broad line region of this quasar has undergone radial velocity variations in a manner resembling Keplerian motion in a binary.
-There certainly are potential caveats, which we discuss in Section 5.2.
+There certainly are potential caveats, which we discuss in Section V.2.
 But the radial velocity curve of J0950 so far is consistent with binary motion, which makes it a strong SBHB candidate.
-We increased the measurement uncertainties of the radial velocity curve to capture the effects of “jitter” from regular quasar variability (guo2019constraining; doan2020improved).
-Quasars can have perturbations or asymmetries within the broad line region that manifest as small, stochastic radial velocity variations, or jitter, occurring over a timescale of the order of a year or less (e.g., barth2015lick).
-This subtle effect is directly measurable in two of J0950’s spectra (see Section 3.2.1).
-doan2020improved empirically determined that jitter can be characterized as fluctuations around a smooth radial velocity curve, distributed approximately according to a Gaussian distribution (see also guo2019constraining).
+We increased the measurement uncertainties of the radial velocity curve to capture the effects of “jitter” from regular quasar variability (Guo et al.
+2019; Doan et al.
+2020).
+Quasars can have perturbations or asymmetries within the broad line region that manifest as small, stochastic radial velocity variations, or jitter, occurring over a timescale of the order of a year or less (e.g., Barth et al.
+2015).
+This subtle effect is directly measurable in two of J0950’s spectra (see Section III.2.1).
+Doan et al.
+(2020) empirically determined that jitter can be characterized as fluctuations around a smooth radial velocity curve, distributed approximately according to a Gaussian distribution (see also Guo et al.
+2019).
 The standard deviation of the distribution is on the scale of a few hundred km s-1.
@@
 The jitter dominates over the measurement errors and is the primary source of uncertainty in the radial velocity at each epoch.
-To fit the radial velocity curve shown in Figure 12, we used radvel (fulton2018radvel), a package for fitting Keplerian orbits to radial velocity time series.
+To fit the radial velocity curve shown in Figure 12, we used radvel (Fulton et al.
+2018), a package for fitting Keplerian orbits to radial velocity time series.
 The best-fitting orbital parameters, determined from the radial velocity curve fit in the observed frame, are provided in Table 4.
@@
 These effects are expected for long period systems.
+In order to improve the sensitivity of the data to longer orbital periods, further monitoring of the radial velocity curve of J0950 is needed for a substantial amount of time.
 The median values of the posterior distributions derived from the MCMC analysis are reported in Table 4.
@@
 d Reduced with 7 degrees of freedom; a small value due to error bars increased by jitter.
-5.4 Exploration of the Supermassive Binary Black Hole Scenario and Discussion
-A lower limit on the total mass of the SBHB in J0950 was initially estimated at M⊙ by runnoe2017large.
+V.4 Exploration of the Supermassive Binary Black Hole Scenario and Discussion
+A lower limit on the total mass of the SBHB in J0950 was initially estimated at M⊙ by Runnoe et al.
+(2017).
 We obtained limits on the mass assuming that the bolometric luminosity does not exceed the Eddington luminosity.
-The bolometric luminosity was estimated by taking the luminosity of J0950’s 2002 spectrum––the brightest state of this object––at 5100 Å, and applying a bolometric correction factor of 10.3 (richards2006spectral).
+The bolometric luminosity was estimated by taking the luminosity of J0950’s 2002 spectrum––the brightest state of this object––at 5100 Å, and applying a bolometric correction factor of 10.3 (Richards et al.
+2006).
 This yields a mass of .
 We estimated another lower mass limit, taking maximum inclination () and a mass ratio of , in the following Keplerian velocity equation,
-This equation was derived under the assumption of an unequal mass binary system, in which the less massive, or secondary, BH is active and its BLR emits the observed H line (see arguments in eracleous2012large; runnoe2015large).
+This equation was derived under the assumption of an unequal mass binary system, in which the less massive, or secondary, BH is active and its BLR emits the observed H line (see arguments in Eracleous et al.
+2012; Runnoe et al.
+2015).
 Using all combinations explored by the MCMC algorithm for the period , the velocity semi-amplitude , and eccentricity , the inferred 68%-confidence lower limit on total mass is and the 99% lower limit is .
@@
 All the lower mass limits above are reasonable in the sense that they are less than , a threshold regarded as unphysical.
-This threshold comes about because the most massive BHs measured are of the order of and because BHs are not expected to accrete material rapidly enough to surpass this mass (e.g., thomas201617; inayoshi2016there; mcconnell2012dynamical).
+This threshold comes about because the most massive BHs measured are of the order of and because BHs are not expected to accrete material rapidly enough to surpass this mass (e.g., Thomas et al.
+2016; Inayoshi and Haiman 2016; McConnell et al.
+2012).
 In line with this argument, the minimum possible inclination for J0950 is .
@@
 To explore that possibility, we estimate the “characteristic radius” of the BLR888This is the length that corresponds to the observed time delay, between fluctuations of continuum and a broad line, .
-using the empirical relation from bentz2013 between the characteristic radius and AGN luminosity at 5100 Å,
+using the empirical relation from Bentz et al.
+(2013) between the characteristic radius and AGN luminosity at 5100 Å,
 with and .
@@
 This value is somewhat higher than our inferred lower limit on the binary semi-major axis of pc.
-The extent of the line-emitting region in the BLR is possibly 45 times greater than (see discussion in runnoe2015large).
+The extent of the line-emitting region in the BLR is possibly 45 times greater than (see discussion in Runnoe et al.
+2015).
 Since we only have a lower limit on the orbital semi-major axis, there appears to be no conflict between this separation and the characteristic radius of the BLR.
@@
 It is therefore critical that we observe this object in the near future, at a rate of at least one spectrum per year to see if this prediction is borne out.
-It is also important to note that our implementation of radial velocity jitter discussed in Section 3.5 is simple, and more sophisticated prescriptions of jitter should be developed and employed.
+It is also important to note that our implementation of radial velocity jitter discussed in Section III.5 is simple, and more sophisticated prescriptions of jitter should be developed and employed.
 The description of jitter would impact the uncertainties on the orbital parameters that are inferred from fitting a radial velocity curve.
 Regardless of whether or not J0950 is a binary, it still displays an exceptional behavior that warrants detailed study.
-5.5 On the Detectability of Doppler Boosting in the Light Curves
-The light curves of J0950 exhibit neither substantial nor periodic variability, as discussed in Section 4.
+V.5 On the Detectability of Doppler Boosting in the Light Curves
+The light curves of J0950 exhibit neither substantial nor periodic variability, as discussed in Section IV.
 In addition to the tests in that section, we phase-folded the light curves using the inferred 33 yr period of J0950, binned the magnitude values by phase, and calculated the median magnitudes and spread within each phase bin.
@@
 Considering a SBHB system, in which a BH carrying the BLR beams light as it orbits, the observed specific flux depends on the angle between the beam’s velocity vector and the observer’s line of sight, through the relativistic Doppler factor , where and .
-For a single source, the observed flux scales as , where is the spectral index (see, for example, dorazio2023review).
+For a single source, the observed flux scales as , where is the spectral index (see, for example, D’Orazio and Charisi, 2023).
 Beaming is most pronounced when the binary is in quadrature.
@@
 Still, the observed variability does not contradict the binary model.
-5.6 Speculation: Evolution of J0950 in the GW landscape
+V.6 Speculation: Evolution of J0950 in the GW landscape
 Given our findings above, it is interesting to speculate about the expected evolution of the GW signal from a system like J0950 in the characteristic strain versus frequency diagram.
@@
 The steps involved in producing the components of the figure are outlined in Appendix C.
-Here, we predict the overall evolution of the system across the PTA and LISA bands using theoretical sensitivity curves by hazboun2019 and robson2019, respectively.
+Here, we predict the overall evolution of the system across the PTA and LISA bands using theoretical sensitivity curves by Hazboun et al.
+(2019) and Robson et al.
+(2019), respectively.
 For the purposes of this illustration we adopt representative values of the mass and mass ratio, as indicated in Figure 13.
@@
 Finally, we estimate the evolutionary time scales for the GW signal from the system, adopting a characteristic binary mass of 10.
-Using the binary decay lifetime from peters1964, it will take 55 Myr for this object to merge.
+Using the binary decay lifetime from Peters (1964), it will take 55 Myr for this object to merge.
 The details of this calculation are given in Appendix C.
@@
 Moreover, if there are other low-redshift systems with properties similar to those adopted above that are now emitting in the PTA band, we will have a chance of detecting them in the near future.
-6 Conclusions
+VI Conclusions
 The spectroscopic supermassive black hole binary candidate J09505128 (abbreviated J0950) was identified in 2012 by the large velocity shift of its broad H line.
@@
 Also presented in both figures are noisy versions of the models.
-These model spectra are referenced in Section 3.2, where we quantify sources of random and systematic errors by cross correlating the broad H emission lines in the full and isolated spectral models.
+These model spectra are referenced in Section III.2, where we quantify sources of random and systematic errors by cross correlating the broad H emission lines in the full and isolated spectral models.
 Appendix B MCMC posterior distributions of orbital parameters from radial velocity fits of J0950
-In Section 5.3, we performed MCMC fits to two radial velocity curves of J0950: one based on the absolute shifts of the isolated broad H models, and another based on the absolute shifts measured from the isolated broad H lines of the data.
+In Section V.3, we performed MCMC fits to two radial velocity curves of J0950: one based on the absolute shifts of the isolated broad H models, and another based on the absolute shifts measured from the isolated broad H lines of the data.
 The fitted orbital parameters are the period (), the time of conjunction (), the natural log of the velocity semi-amplitude (ln ), and and .
@@
 Appendix C Evolution of a system like J0950 in the PTA and LISA gravitational-wave landscape
-The characteristic strain in the detector frame of an eccentric binary at the th harmonic, can be determined using a simplified expression from breivik2020,
-Here, is the chirp mass, is luminosity distance, is the detector frame GW frequency, is an eccentricity enhancement factor, and is a function involving a series of Bessel functions of the first kind, that captures the contribution of each harmonic and eccentricity to the gravitational radiation power spectrum (see peters&mathews1963).
+The characteristic strain in the detector frame of an eccentric binary at the th harmonic, can be determined using a simplified expression from Breivik et al. 2020,
+Here, is the chirp mass, is luminosity distance, is the detector frame GW frequency, is an eccentricity enhancement factor, and is a function involving a series of Bessel functions of the first kind, that captures the contribution of each harmonic and eccentricity to the gravitational radiation power spectrum (see Peters and Mathews 1963).
 To produce the evolutionary tracks of a system like J0950 for the different harmonics shown in Figure 13, we considered a range of eccentricities from 0.65 to approximately zero.
@@
 The system was then evolved from the frequency where the eccentricity is approximately zero to the frequency at merger, maintaining a fixed .
-Recall that for this exercise, we defined “merger” as the separation at ISCO in Section 5.6.
-The timescales provided in Section 5.6 for such a system to merge and its GW signal to be detected by PTA experiments and LISA, were computed using the decay lifetime equation given by peters1964,
+Recall that for this exercise, we defined “merger” as the separation at ISCO in Section V.6.
+The timescales provided in Section V.6 for such a system to merge and its GW signal to be detected by PTA experiments and LISA, were computed using the decay lifetime equation given by Peters 1964,
 where .